### v2

# Structured coalescents, coagulation equations and multi-type branching processes

Abstract.
Consider a structured population consisting of $d$ colonies, with migration rates that are proportional to a positive parameter $K$. We sample $N_{K}$ individuals distributed evenly across the $d$ colonies and trace their ancestral lineages back. Within a colony, we assume that pairs of ancestral lineages coalesce at a constant rate, as in a Kingman’s coalescent. We identify each ancestral lineage with the set, or block, of its descendants in the sample and we encode the state of the system using a $d$-dimensional vector of empirical measures; the $i$-th component records the blocks present in colony $i$ and the initial location of the lineages composing each block. We are interested in the asymptotic behaviour of the process of empirical measures as $K\to\infty$. We consider two regimes: the critical sampling regime, where $N_{K}\sim K$ and the large sampling regime where $N_{K}\gg K$. After an appropriate time-space scaling, we show that the process of empirical measures converges to the solution of a $d$-dimensional coagulation equation. In the critical sampling regime, the solution can be represented in terms of a multi-type branching process. In the large sampling regime, the solution can be represented in terms of the entrance law of a multi-type continuous state branching process.
MSC 2020. Primary: 60J90; 60G57 Secondary: 60J95; 60B10; 60J80; 35Q92
Keywords. Structured coalescent; coagulation equations; multi-type branching processes; multi-type Feller-diffusions
1. Introduction
Coalescents, branching processes, and coagulation equations represent three fundamental approaches for modeling the dynamics of general interacting particle systems. Each of these captures a distinct yet interconnected facet of stochastic evolution: coalescents describe the merging of ancestral lineages backward in time [25, 24, 31, 38, 40]; branching processes model population growth and reproduction mechanisms forward in time [19, 3, 20, 27, 29, 15]; and coagulation equations provide deterministic approximations for the evolution of cluster sizes in systems undergoing mass-conserving mergers [26, 13, 12, 35].
The relationship between these objects is theoretically rich and practically significant. For example, it is well known [2, 11, 7] that the mean field limit of the, properly scaled, evolution of block sizes in Kingman’s coalescent is given by a Smoluchowski’s coagulation. Furthermore, probabilistic representations of one-dimensional coagulation equations and single-type branching processes can be found, for example, in [11]. In the multi-type setting these connections are more involved. For instance, in the recent works [22, 17] it has been shown that the genealogies of multi-type continuous state branching processes can be described (at least locally) by means of exchangeable multi-type coalescents [23, 32, 14]. Along the same lines, one of the present authors (E.S.) introduced a nested coalescent model [28, 9], where gene lineages evolve within a larger species tree. That study established a connection between the nested coalescent and a transport-coagulation equation, and demonstrated that the corresponding deterministic PDE admits a stochastic representation in terms of a branching CSBP. In this work, we establish a similar connection between a structured Kingman coalescent [36], a multidimensional Smoluchowski coagulation equation, and multi-type branching processes. This generalizes some of the aforementioned results available for the Kingman coalescent and the nested coalescent.
Our starting point is the classical structured coalescent of [36] where a population is structured into $d$ colonies connected through migration. We sample $N_{K}$ individuals, distributed across the $d$ colonies, and trace their genealogies backwards in time. Within a colony, pairs of ancestral lineages are assumed to coalesce as in a Kingman coalescent.
We will examine the regime of fast migration, where migration occurs at a much faster rate than coalescence. Specifically, we consider migration rates of order $\mathcal{O}(K)$ with $K\to\infty$. This regime has been widely investigated in the literature in the context of population genetics (see, e.g., [10] for a review). For a fixed sample size, Norborg and Krone demonstrated in [33, 34] that the genealogical structure effectively collapses in the fast migration limit, resulting in an averaged behavior across colonies. In this extreme case of complete structural collapse, the process converges to the standard one-dimensional Kingman coalescent, which corresponds to a fully mixed population.
A key assumption in the previous averaging result is that the sample size remains fixed as the migration scale $K$ increases. In contrast, this article focuses on a different asymptotic regime, where the sample size $N_{K}$ grows with the migration scale, i.e., $N_{K}\to\infty$ as $K\to\infty$. More precisely, we investigate two distinct sampling regimes: the critical sampling regime, where $N_{K}\sim K$, and the large sampling regime, where $N_{K}\gg K$. In both cases, each ancestral lineage retains information about the block of sampled individuals it traces back to and the aforementioned averaging does not hold anymore. To capture this result, we represent the system’s state as a $d$-dimensional vector of empirical measures, where the $i$-th component captures the blocks present in colony $i$ and their corresponding block configurations—the initial locations of the lineages within each block. We will show that this process converges to the solution of a multi-dimensional Smoluchowski-type coagulation equation under an appropriate small time scaling. Further, we will show that the resulting equation admits a natural probabilistic interpretation in terms of multi-type branching processes. Taken together, our approach hints at a unified framework that connects structured coalescents, coagulation equations, and multi-type branching processes.
A natural question arising from this work is how to characterize the Site Frequency Spectrum (SFS) of the structured coalescent in the regime of fast migration and large sample sizes. Assuming that neutral mutations occur at a constant rate along the branches of the coalescent tree, the $i^{\text{th}}$ component of the SFS (for $i\in[N_{K}-1]$) represents the number of mutations present in exactly $i$ leaves. This question has been studied in the context of $\Lambda$- and $\Xi$-coalescents (see, e.g., [4, 18]), where it has been shown that, in the limit of large sample size, the lower end of the SFS is influenced by the small-time behavior of the coalescent process. We believe that the small-time asymptotics derived in our work may yield analogous results in the structured coalescent setting. Investigating this will be the focus of future research.
The remainder of the paper is organized as follows. Section 2 introduces the model and states the main results. Subsequent sections are devoted to proving these results. Section 3 proves the convergence of the generators of the empirical measure process and states a comparison result that allows us to bound the number of blocks in our structured coalescent between the number of blocks of in two Kingman coalescents with different coalescent rates. Section 4 establishes the tightness of the sequence of empirical measures and characterizes their accumulation points. We conclude the proofs of our main results in Section 5.
2. Model and main results
In this section we formalize the definition of the structured coalescent that will be the object of our analysis and we state our main results.
2.1. The model
We consider a prototype model for the genealogies of a large population divided into $d$ colonies, where $d\in{\mathbb{N}}$ will remain fixed throughout the manuscript. The model takes the form of a structured coalescent that evolves according to the following principles. Within a colony, blocks coalesce as in Kingman’s coalescent (i.e. at a constant rate per pair of blocks), while blocks can also migrate between colonies at rates proportional to a scaling parameter $K$. The aim of this article is to analyze the effect of sample size $N_{K}$ in the corresponding ancestral structures, at small times, as $K\to\infty$.
Let us now formalize the previous description. Each individual in the population is identified by a unique number in $[N_{K}]\coloneqq\{1,\ldots,N_{K}\}$; each colony is identified by a unique number in $[d]$, which we will often refer to as a color. The state of the system is encoded by a structured partition as defined below.
Definition 2.1 (Structured/colored partition).
Let $N\in\mathbb{N}$. We refer to the elements of ×[d]$\mathbb{N}\times[d]$ as colored (positive) integers. We say that $S\subset[N]\times[d]$ is a coloring of $[N]$ if $S=\{(1,c_{1}),\ldots,(N,c_{N})\}$ for some $c_{1},\ldots,c_{N}\in[d]$. We refer to ${\boldsymbol{\pi}}\coloneqq(\pi_{i})_{i\in[d]}$ as a colored partition of a coloring $S$ of $[N]$ if $\pi_{1},\ldots,\pi_{d}$ are disjoint (possibly empty) collections of non-empty subsets of $S$ and $\pi_{1}\cup\pi_{2}\cup\ldots\cup\pi_{d}$ is a partition of $S$; we call the elements of $\pi_{i}$ blocks of color $i$. If $S$ is a coloring of $[N]$, we denote by ${\mathcal{S}}_{S}^{N}$ the set of all colored partitions of $S$ and by ${\mathcal{S}}^{N}$ the set of all colored partitions of some coloring of $[N]$. We equip ${\mathcal{S}}_{S}^{N}$ and ${\mathcal{S}}^{N}$ with the discrete topology.
Example 2.2.
$S={\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}1},{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}2},{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}3,4},{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}5}$
is a coloring of $[N=5]$ with $d=2$ colors.
${{\boldsymbol{\pi}}}=\underbrace{{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\{}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}1}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}5}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\}}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\{}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}3}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\}}}_{\pi_{1}}\underbrace{{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}\{}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}2}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}4}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}\}}}_{\pi_{2}}$
is a colored partition of $S$ with $2$ blue blocks and 1 green block.
Now, we describe the anticipated multi-type coalescent as a Markov process on the set of colored partitions. Fix $K\geq 1$ and $N_{K}\in\mathbb{N}$. Let ${\boldsymbol{W}}\coloneqq(w_{i,j})_{i,j\in[d]}$ be a primitive matrix and ${\boldsymbol{\alpha}}\coloneqq(\alpha_{i})_{i\in[d]}\in{\mathbb{R}}_{+}^{d}$. The multi-type coalescent ${\boldsymbol{\Pi}}^{K}\coloneqq(\mathbf{\Pi}^{K}(t))_{t\geq 0}$ is the continuous-time Markov chain on ${\mathcal{S}}^{N_{K}}$ which is characterized by the following dynamics:
  1. (1)
    At time 00, start with a coloring of $[N_{K}]$. The partition $\Pi_{0}$ is the set of singletons whose coloring coincides with the coloring of $[N_{K}]$. For instance, for the coloring of $S$ in Example 2.2,
    $\Pi^{K}_{0}={\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\{}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}1}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\}}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}\{}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}2}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}\}}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\{}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}3}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\}}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\{}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}4}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\}}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}\{}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}5}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}\}}.$
  2. (2)
    Each block changes its color from $i$ to $j$ at rate $Kw_{i,j}$. For instance
    {15}{15}at rate ${\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\{}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}1}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}5}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\}}\to{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}\{}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}1}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}5}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}\}}\ \ \mbox{at rate $Kw_{1,2}$.}$
  3. (3)
    For every $i$ in $[d]$, each pair of blocks of color $i$ coalesce into a single block of color $i$ at rate $\alpha_{i}$. For instance
[…]
    Pairs of different colors do not coalesce.
If we envision the coalescent as a random ultrametric tree, the color of a block and its internal coloring can be interpreted as follows. The color of a block at time $t$ is the position (colony) of the corresponding lineage. The internal coloring provides the labels and colors of the leaves supported by this lineage $t$ units of time in the past.
Assumption 2.3.
Let $L_{i}^{K}(t)$ be the number of blocks of color $i$ at time $t$. There is a vector ${\boldsymbol{\beta}}\coloneq(\beta_{i})_{i\in[d]}\in{\mathbb{R}}_{+}^{d}$, such that, for each $i\in[d]$,
$\frac{L_{i}^{K}(0)}{N_{K}}\xrightarrow[K\to\infty]{}\beta_{i}.$
Since the matrix ${\boldsymbol{W}}$ is primitive, it follows that it admits a unique stationary distribution, which we denote by ${\boldsymbol{\xi}}\coloneqq(\xi_{i})_{i\in[d]}$.
2.2. The empirical measure
In this work we will focus on the asymptotic properties of a functional of the coalescent $\boldsymbol{\Pi}^{K}$, which will code for the colony sizes, but also for the color-configurations within blocks (see Fig. 1). To encode this information, we will need a few definitions. We will say that the configuration of a block is ${{\boldsymbol{k}}}=(k_{i})_{i\in[d]}\in[N_{K}]_{0}^{d}$ if and only if the block contains $k_{1}$ elements of color $1$, $k_{2}$ elements of color $2$ etc. For instance, the configuration of the blue block ${\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\{}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}1}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}3}{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}5}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}\}}$ is ${\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4}(1},{\color[rgb]{0.475,0.7,0.55}\definecolor[named]{pgfstrokecolor}{rgb}{0.475,0.7,0.55}2}{\color[rgb]{0.0,0.14,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.14,0.4})}$.
Figure 1. An illustration of the multi-type coalescent for $d=3$, $N_{K}=10$. Blocks are classified according to their colors, which code for the different colonies. A change in color represents a migration event. Specifically, multi-colored squares indicate a migration from colony $i$ to colony $j$, where the upper color corresponds to the origin colony $i$, and the lower color to the destination colony $j$.
For every time $t\in\mathbb{R}_{+},i\in[d]$, we define the empirical measure $\nu_{i}^{K}(t,\cdot)\in{\mathcal{M}}_{f}([N_{K}]^{d}_{0})$ via
$\ \nu_{i}(t,\{{\boldsymbol{k}}\})\ \coloneqq\ \#\{\mbox{blocks of color $i$ with configuration ${{\boldsymbol{k}}}$}\},\quad{{\boldsymbol{k}}}=(k_{i})_{i\in[d]}\in[N_{K}]_{0}^{d}.$
In words, $\nu_{i}(t,\{{\boldsymbol{k}}\})$ is the number of lineages at time $t$ located in colony $i$ carrying $k_{1}$ leaves of color $1$ etc. As anticipated at the beginning of the section, our aim is to understand the interplay between the sampling size $N_{K}$ and the scale $K$ at which migrations (change of color) occur in the model. In our analysis, we will therefore distinguish between the two regimes depending on the asymptotic behavior of $\gamma_{K}\coloneqq N_{K}/K$. More precisely, we consider
  1. (1)
    The critical sampling regime:
    $\gamma_{K}\underset{K\to\infty}{\longrightarrow}c>0.$
  2. (2)
    The large sampling regime:
    $\gamma_{K}\underset{K\to\infty}{\longrightarrow}\infty.$
In the critical sampling regime, the rates of migration and coalescence events will both be of the same order as long as the total number of blocks remains of order $\mathcal{O}(K)$. In contrast, in the large sampling regime, coalescence events will dominate the dynamics as long as the total number of blocks is of a higher order than $K$, and only when the total number of blocks reaches order $K$ will coalescence events and migrations synchronize on the same time scale. We will therefore slow down time by $1/K$, so that the time frame in which coalescence dominates the dynamics is very short, and hence the strengths of migration and coalescence equilibrate very quickly (see Fig. 2 for an illustration of the time scaling). As a consequence, in the large sampling regime, blocks will quickly become very large, and we will therefore need to introduce an appropriate space scaling. More precisely, and to avoid unnecessary case distinctions when scaling blocks, we introduce the space-scaling parameter
$s_{K}\coloneq\begin{cases}1,&\quad\text{in the critical sampling}\\ \gamma_{K},&\quad\text{in the large sampling}.\end{cases}$ (2.1)
Note that in both regimes
$b\coloneqq\sup_{K}\gamma_{K}/s_{K}<\infty.$ (2.2)
With this intuition in mind, we consider the rescaled measure valued process $(\mu^{K}_{i}(t,{\rm d}x))_{t\geq 0}$ valued in ${\mathcal{M}}_{f}(\mathbb{R}_{+}^{d})$ equipped with weak topology as
$\ \ \ \mu_{i}^{K}(t,{\rm d}x)\ \coloneqq\frac{1}{K}\ \nu^{K}_{i}\left(\frac{t}{K},s_{K}\,{\rm d}x\right),\qquad i\in[d],t\geq 0.$
Figure 2. An illustration of the effect of the time scaling $t\mapsto t/K$.
2.3. Convergence
In this section we state the two main results concerning the convergence of the empirical measures $(\mu_{i}^{K})_{i\in[d]}$ as $K\to\infty$. Let us anticipate that our results will differ substantially depending on the parameter regime under consideration.
We start by introducing some notation. If $\eta_{1},\eta_{2}\in{\mathcal{M}}_{f}({\mathbb{R}}^{d})$ we denote by $\eta_{1}\star\eta_{2}$ their convolution, i.e.
$\eta_{1}\star\eta_{2}(B)=\int\mathds{1}_{B}({\boldsymbol{x}}+{\boldsymbol{y}})\eta_{1}(\mathrm{d}{\boldsymbol{x}})\eta_{2}(\mathrm{d}{\boldsymbol{y}}).$
In addition, if $\eta\in{\mathcal{M}}_{f}({\mathbb{R}}^{d})$ and $f:{\mathbb{R}}^{d}\to{\mathbb{R}}$ is integrable with respect to $\eta$, we set
$\langle\eta,f\rangle\coloneqq\int f({\boldsymbol{x}})\eta(\mathrm{d}{\boldsymbol{x}}).$
If the involved measures have discrete support, the integrals can, as usual, be understood as sums.
Now we state our main convergence result in the critical regime.
Theorem 2.4 (Critical sampling).
Assume that $\gamma_{K}\to{c}$ as $K\to\infty$. If Assumption 2.3 holds, then $(\mu_{i}^{K})_{i\in[d]}$ weakly converges, as $K\to\infty$, to the solution of the $d$-dimensional discrete coagulation equation
$\partial_{t}u_{i}(t,{\boldsymbol{n}})$ $=\alpha_{i}\left(\frac{1}{2}u_{i}\star u_{i}(t,{\boldsymbol{n}})-\left<u_{i}(t),1\right>u_{i}(t,{\boldsymbol{n}})\right)+\sum_{j\in[d]\setminus\{i\}}w_{j,i}u_{j}(t,{\boldsymbol{n}})-w_{i,j}u_{i}(t,{\boldsymbol{n}}),$ (2.3)
$u_{i}(0,{\boldsymbol{n}})$ $=c\beta_{i}\delta_{{\boldsymbol{e}}_{i},{\boldsymbol{n}}},$ (2.4)
for all $t\geq 0$, ${\boldsymbol{n}}\in\mathbb{N}_{0}^{d}\setminus\{\boldsymbol{0}\}$ and $i\in[d]$, where $({\boldsymbol{e}}_{i})_{i\in[d]}$ denotes the canonical basis of $\mathbb{R}^{d}$.
The dynamics of the coagulation equation is very intuitive. When two blocks coalesce, the resulting block configuration is the sum of the block configurations of the coalescing blocks; hence the convolution. Moreover, whenever a block with a given configuration merges with another block, we lose that configuration. The remaining sum is due to migration: we gain a block configuration in colony $i$ whenever a block with that configuration migrates from another colony to colony $i$; we lose a block configuration whenever a block with that configuration migrates from colony $i$ to another colony $i$. The initial condition is due to the fact that at time $t=0$ there are approximately $\beta_{i}K$ singletons in colony $i$.
The next theorem covers the large sampling case.
Theorem 2.5 (Large sampling).
Assume that $\gamma_{K}\to\infty$ as $K\to\infty$. If Assumption 2.3 holds, then $(\mu_{i}^{K})_{i\in[d]}$ weakly converges, as $K\to\infty$, to the weak solution of the $d$-dimensional continuous coagulation equation
$\partial_{t}u_{i}=\alpha_{i}\left(\frac{1}{2}u_{i}\star u_{i}-u_{i}\left<u_{i},1\right>\right)\ +\ \sum_{j\in[d]\setminus\{i\}}w_{ji}u_{j}-w_{ij}u_{i},\quad i\in[d],\,t>0$ (2.5)
$\lim_{t\to 0}\int\left(1-e^{-\left<\boldsymbol{\lambda},{\boldsymbol{x}}\right>}\right)u_{i}(t,{\rm d}{\boldsymbol{x}})=\lambda_{i}\beta_{i},\quad\boldsymbol{\lambda}=(\lambda_{i})_{i\in[d]}>\boldsymbol{0},\,i\in[d].$ (2.6)
Remark 2.6.
By weak solution, we mean that for $f\in C_{b}(\mathbb{R}_{+}^{d})$
$\langle u_{i}(t),f\rangle=\ \langle u_{i}(0),f\rangle$ $+\frac{\alpha_{i}}{2}\int_{0}^{t}(\langle(u_{i}\star u_{i})(s),f\rangle-2\ \langle u_{i}(s),1\rangle\ \langle u_{i}(s),f\rangle)\mathrm{d}s$
$+\sum_{j\in[d]\setminus\{i\}}\int_{0}^{t}(w_{j,i}\ \langle u_{j}(s),f\rangle-w_{i,j}\text{ }\langle u_{i}(s),f\rangle)\mathrm{d}s.$
2.4. Stochastic representation
In this section we provide stochastic representations to the coagulation equations obtained in the previous section. We provide here the derivation only on the large sampling case; the critical sampling case will be treated analogously in Appendix A. Therefore, throughout this section, we assume that $\gamma_{K}\to\infty$ as $K\to\infty$.
Let ${{\boldsymbol{u}}}$ be the solution of the continuum coagulation equation (2.5) under (2.6). Define
$v_{i}(t,\boldsymbol{\lambda})\ \coloneqq\frac{1}{\beta_{i}}\int_{\mathbb{R}_{+}^{d}}\left(1-e^{-\left<{{\boldsymbol{\lambda}}},{{\boldsymbol{x}}}\right>}\right)u(t,{\rm d}{\boldsymbol{x}}).$
A direct computation shows that ${\boldsymbol{v}}$ solves the multi-dimensional ODE
$\partial_{t}v_{i}(t,\boldsymbol{\lambda})\ =\ -\psi_{i}({{\boldsymbol{v}}}(t,\boldsymbol{\lambda}))\ \mbox{and $v_{i}(0,\boldsymbol{\lambda})=\lambda_{i},$}$
where
$\psi_{i}({\boldsymbol{\lambda}})\ =\ \frac{1}{2}(\alpha_{i}\beta_{i})\lambda_{i}^{2}-\sum_{j\in[d]\setminus\{i\}}\bigg{(}w_{j,i}\frac{\beta_{j}}{\beta_{i}}\lambda_{j}-w_{i,j}\lambda_{i}\bigg{)}.$
We identify ${\boldsymbol{v}}$ as the Laplace exponent of the ${\boldsymbol{\psi}}$-CSBP, or equivalently the $d$-dimensional Feller diffusion ${{\boldsymbol{Z}}}$ solution of the SDE (see e.g. [37, Chap. 4.2, Eq. (4.4)])
${\rm d}Z_{i}(t)\ =\ \sqrt{\alpha_{i}\beta_{i}Z_{i}(t)}{\rm d}B_{i}(t)\ +\ \sum_{j\in[d]\setminus\{i\}}\Big{(}w_{j,i}\frac{\beta_{j}}{\beta_{i}}Z_{j}(t)-w_{i,j}Z_{i}(t)\Big{)}{\rm d}t,\quad i\in[d],$ (2.7)
where ${\boldsymbol{B}}$ is a standard $d$-dimensional Brownian motion. Thus, for every ${\boldsymbol{x}}=(x_{i})_{i\in d]}$, we have
$\mathbb{E}_{{\boldsymbol{x}}}(e^{-{\boldsymbol{\lambda}}\cdot{\boldsymbol{Z}}_{t}})\ =\ e^{-\left<{{\boldsymbol{x}}},{\boldsymbol{v}}(t,{\boldsymbol{\lambda}})\right>}.$ (2.8)
Let $Q_{i}$ be the entrance law of the process from state $i$, i.e.
$\ \ Q_{i}(t,A)\ \coloneqq\ \lim_{x\to 0+}\frac{\mathbb{P}_{x{\boldsymbol{e}}_{i}}(Z_{t}\in A)}{x},\quad t>0.$
Using standard arguments (see e.g. [30, Thm. 3.13]), one can deduce from (2.8) that for $t>0$
$v_{i}(t,\boldsymbol{\lambda})\ =\ \int_{\mathbb{R}_{+}^{d}}\left(1-e^{-\left<{{\boldsymbol{\lambda}}},{{\boldsymbol{x}}}\right>)}\right)Q_{i}(t,\mathrm{d}{\boldsymbol{x}}).$
But recall that
$v_{i}(t,\boldsymbol{\lambda})=\frac{1}{\beta_{i}}\int_{\mathbb{R}_{+}^{d}}\left(1-e^{-\left<{{\boldsymbol{\lambda}}},{{\boldsymbol{x}}}\right>}\right)u_{i}(t,{\rm d}{\boldsymbol{x}}).$
We can then invert the Laplace transform and get the following result.
Theorem 2.7.
The solution of the coagulation equation (2.5) with initial condition (2.6) admits the stochastic representation
$u_{i}(t,{{\rm d}{\boldsymbol{x}}})\ =\ \beta_{i}Q_{i}(t,{{\rm d}{\boldsymbol{x}}}),\quad t>0,\,i\in[d],$
where $Q_{i}$ is the entrance law of ${{\boldsymbol{Z}}}$ of the $d$-dimensional Feller diffusion (2.7) starting from state $i$. In particular, the solution to (2.5) with initial condition (2.6) is unique.
For the critical sampling, we will prove the following analogous result in Appendix A.
Theorem 2.8.
Assume that for each $i\in[d]$, we have
$d_{i}\coloneqq\frac{c\alpha_{i}\beta_{i}}{2}-\sum_{j\in[d]\setminus\{i\}}\left(\frac{\beta_{j}}{\beta_{i}}w_{j,i}-w_{i,j}\right)>0.$
Then, the solution of the coagulation equation (2.3) with initial condition (2.4) admits the stochastic representation
$u_{i}(t,{\boldsymbol{n}})\ =\ {c}\beta_{i}{\mathbb{P}}_{{\boldsymbol{e}}_{i}}\big{(}{\boldsymbol{Z}}(t)\ =\ {\boldsymbol{n}}\big{)},\quad t\geq 0,\,{\boldsymbol{n}}\in\mathbb{N}_{0}^{d}\setminus\{\boldsymbol{0}\},\,i\in[d],$
where ${\boldsymbol{Z}}(t)=(Z_{i}(t))_{i\in[d]}$ denotes a continuous-time multi-type branching process, such that a particle of color $i\in[d]$
  • branches at rate $\frac{{c}\alpha_{i}\beta_{i}}{2}$,
  • dies at rate $d_{i}$,
  • makes a transition from $i$ to $j$ at rate $\frac{\beta_{j}}{\beta_{i}}w_{j,i}$.
Remark 2.9.
In particular, the previous result holds for sufficiently large values of $c$. This aligns with the large sampling case, in which $c$ can be understood as being infinite.
The uniqueness of the solution to the coagulation equation (2.3) with initial condition (2.4) will be addressed in the Appendix; the assumption $d_{i}\geq 0$ will be not required.
2.5. Proof strategy
The remainder of the manuscript is devoted to the proofs of Theorems 2.4 and 2.5. In Section 3.1, we characterize the generator of the process of empirical measures and establish its convergence in a precise sense. In Section 3.4 we establish a coupling allowing us to compare the number of blocks in our multi-type coalescent to the number of blocks in Kingman’s coalescent; the resulting bounds will play an important role in later sections. In Section 4, we prove the tightness of the sequence of empirical measures and characterize its accumulation points. Finally, in Section 5, we complete the proofs of Theorems 2.4 and 2.5, offering further insight into the role of the initial condition in the large sampling regime (i.e. Eq. (2.4)).
3. Convergence of the generators and Kingman bounds
In this section we describe the action of the infinitesimal generator $A^{K}$ of the process of empirical measures ${\boldsymbol{\mu}}^{K}\coloneqq(\mu_{1}(t,\cdot),\ldots\mu_{d}(t,\cdot))_{t\geq 0}$ on a special class of functions. We will use $A^{K}$ to derive properties of the process ${\boldsymbol{\mu}}^{K}$. In the second half of this section we will prove the convergence as $K\to\infty$ of the generators $A^{K}$, in an appropriate sense.
3.1. The action of the generator
Let ${\mathcal{H}}$ be the set of real-valued functions on $(\mathcal{M}_{f}(\mathbb{R}_{+}^{d}))^{d}$ of the form
${\boldsymbol{p}}\coloneqq(p_{i})_{i\in[d]}\in(\mathcal{M}_{f}(\mathbb{R}_{+}^{d}))^{d}\mapsto H^{F,{\boldsymbol{f}}}({\boldsymbol{p}})\coloneqq F(\langle{\boldsymbol{p}},{\boldsymbol{f}}\rangle)$ , (3.1)
for some $F\in C({\mathbb{R}}^{d})$ and ${\boldsymbol{f}}\coloneqq(f_{i})_{i\in[d]}$ with $f_{i}\in C(\mathbb{R}_{+}^{d})$ for $i\in[d]$, where
$\langle{\boldsymbol{p}},{\boldsymbol{f}}\rangle\coloneqq(\langle p_{i},f_{i}\rangle)_{i\in[d]}.$
To describe the action of the generator $A^{K}$ of the process of empirical measures ${\boldsymbol{\mu}}^{K}$ on functions in ${\mathcal{H}}$, we write
$A^{K}H^{F,{\boldsymbol{f}}}({\boldsymbol{p}})=A^{K}_{M}H^{F,{\boldsymbol{f}}}({\boldsymbol{p}})+A^{K}_{C}H^{F,{\boldsymbol{f}}}({\boldsymbol{p}}),$
where the operators $A^{K}_{M}$ and $A_{C}^{m}$ account for the contributions of migrations and coalescences, respectively. For the migration part, recall that any block-configuration ${\boldsymbol{c}}$ present in colony $i$ migrates to colony $j$ at rate $w_{i,j}K$, which has the effect on ${\boldsymbol{\mu}}^{K}$ of adding to coordinate $j$ a mass $1/K$ at ${\boldsymbol{c}}/s_{K}$ and removing from coordinate $i$ a mass $1/K$ at ${\boldsymbol{c}}/s_{K}$ (see Fig. 3 (right)). Since we are scaling time by $1/K$, we obtain
$A^{K}_{M}H^{F,{\boldsymbol{f}}}({\boldsymbol{p}})=K\sum_{i,j\in[d]}w_{i,j}\sum_{{\boldsymbol{c}}\in\mathbb{N}_{0}^{d}}p_{i}\Big{(}\Big{\{}\frac{{\boldsymbol{c}}}{s_{K}}\Big{\}}\Big{)}\bigg{[}F\bigg{(}\langle{\boldsymbol{p}},{\boldsymbol{f}}\rangle+\frac{\Delta_{i,j}{\boldsymbol{f}}(\frac{{\boldsymbol{c}}}{s_{K}})}{K}\bigg{)}-F\left(\langle{\boldsymbol{p}},{\boldsymbol{f}}\rangle\right)\bigg{]},$ (3.2)
where $\Delta_{i,j}{\boldsymbol{f}}={\boldsymbol{e}}_{j}f_{j}-{\boldsymbol{e}}_{i}f_{i}$.
Figure 3. An illustration of the two types of transitions. Left: coalescence (the block color remains the same while the resulting configuration is the sum of the configurations of the two merging blocks). Right: migration (the block color changes while the configuration remains the same).
For the coalescent part, recall that any pair of block configurations ${\boldsymbol{c}}_{1}$, ${\boldsymbol{c}}_{2}$ present in colony $i$ coalesces at rate $\alpha_{i}$, which has the effect on ${\boldsymbol{\mu}}^{K}$ of adding to the coordinate $i$ a mass $1/K$ at $({\boldsymbol{c}}_{1}+{\boldsymbol{c}}_{2})/s_{K}$ and removing from the coordinate $i$ a mass $1/K$ at ${\boldsymbol{c}}_{1}/s_{K}$ and ${\boldsymbol{c}}_{2}/s_{K}$ (see Fig. 3 (left)). Since we are scaling time by $1/K$, if we distinguish between the cases where ${\boldsymbol{c}}_{1}\neq{\boldsymbol{c}}_{2}$ and ${\boldsymbol{c}}_{1}={\boldsymbol{c}}_{2}$, we obtain
$A^{K}_{C}H^{F,{\boldsymbol{f}}}({\boldsymbol{p}})=\frac{K}{2}\sum_{i\in[d]}{\alpha_{i}}\sum_{{\boldsymbol{c}}_{1}\neq{\boldsymbol{c}}_{2}}p_{i}\Big{(}\Big{\{}\frac{{\boldsymbol{c}}_{1}}{s_{K}}\Big{\}}\Big{)}p_{i}\Big{(}\Big{\{}\frac{{\boldsymbol{c}}_{2}}{s_{K}}\Big{\}}\Big{)}\bigg{[}F\bigg{(}\langle{\boldsymbol{p}},{\boldsymbol{f}}\rangle+\frac{\Delta_{i}f(\frac{({\boldsymbol{c}}_{1},{\boldsymbol{c}}_{2})}{s_{K}})}{K}\bigg{)}-F\left(\langle{\boldsymbol{p}},{\boldsymbol{f}}\rangle\right)\bigg{]}$
$+\frac{1}{2}\sum_{i\in[d]}\alpha_{i}\sum_{{\boldsymbol{c}}\in\mathbb{N}_{0}^{d}}p_{i}\Big{(}\Big{\{}\frac{{\boldsymbol{c}}}{s_{K}}\Big{\}}\Big{)}\Big{(}Kp_{i}\Big{(}\Big{\{}\frac{{\boldsymbol{c}}}{s_{K}}\Big{\}}\Big{)}-1\Big{)}\bigg{[}F\bigg{(}\langle{\boldsymbol{p}},{\boldsymbol{f}}\rangle+\frac{\Delta_{i}f(\frac{({\boldsymbol{c}},{\boldsymbol{c}})}{s_{K}})}{K}\bigg{)}-F\left(\langle{\boldsymbol{p}},{\boldsymbol{f}}\rangle\right)\bigg{]},$ (3.3)
where $\Delta_{i}{\boldsymbol{f}}({\boldsymbol{x}}_{1},{\boldsymbol{x}}_{2})={\boldsymbol{e}}_{i}(f_{i}({\boldsymbol{x}}_{1}+{\boldsymbol{x}}_{2})-f_{i}({\boldsymbol{x}}_{1})-f_{i}({\boldsymbol{x}}_{2}))$.
3.2. Useful bounds
The next result provides bounds for some functionals of the measures $(\mu_{i}^{K})_{i\in[d]}$, which will be used in Section 4 (see proof of Lemma 4.8) and Section 5 (see proof of Proposition 5.1).
Lemma 3.1.
For any $t\geq 0$, we have
$\mathbb{E}\bigg{[}\sum_{i\in[d]}\langle\mu_{i}^{K}(t),\|\cdot\|_{1}^{2}\rangle\bigg{]}\leq b^{2}\Big{(}\frac{1}{\gamma_{K}}+\alpha_{\mathrm{max}}t\Big{)}$ (3.4)
and
$\mathbb{E}\Bigg{[}\bigg{(}\sum_{i\in[d]}\langle\mu_{i}^{K}(t),\|\cdot\|_{1}^{2}\rangle\bigg{)}^{2}\Bigg{]}\leq e^{\frac{2\alpha_{\mathrm{max}}}{K}t}b^{4}\Big{(}\frac{1}{\gamma_{K}^{2}}+\frac{2\alpha_{\mathrm{max}}}{\gamma_{K}}t+\alpha_{\mathrm{max}}^{2}t^{2}\Big{)},$ (3.5)
where $\alpha_{\mathrm{max}}\coloneqq\max_{i\in[d]}\alpha_{i}$ and $b$ is defined in (2.2).
Proof.
We begin with the proof of (3.4). Note that for $F_{1}({\boldsymbol{x}})\coloneqq\sum_{i\in[d]}x_{i}$ and $f({\boldsymbol{x}})\coloneqq\|{\boldsymbol{x}}\|_{1}^{2}$, we have
$H^{F_{1},{\boldsymbol{f}}}({\boldsymbol{p}})=\sum_{i\in[d]}\langle{\boldsymbol{p}},\|\cdot\|_{1}^{2}\rangle.$
Hence $H^{F,{\boldsymbol{f}}}\in{\mathcal{H}}$. Moreover, Eq. (3.2) yields $A^{K}_{M}H^{F_{1},{\boldsymbol{f}}}(\boldsymbol{{\boldsymbol{p}}})=0$. To deal with coalescences, we group block-configurations in sets of the form $C_{\ell}^{K}\coloneqq\{{\boldsymbol{c}}/s_{K}:\lVert{\boldsymbol{c}}\rVert_{1}=\ell\}$, and we use Eq. (3.1) together with the inequality $\sum_{i}|a_{i}b_{i}|\leq(\sum_{i}|a_{i}|)(\sum_{i}|b_{i}|)$, to obtain
$A^{K}_{C}H^{F_{1},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K}(t))$ $\leq\frac{\alpha_{\mathrm{max}}}{2}\sum_{\ell_{1},\ell_{2}\in\mathbb{N}}\bigg{(}\sum_{i\in[d]}\mu_{i}^{K}(t,C_{\ell_{1}}^{K})\bigg{)}\bigg{(}\sum_{i\in[d]}\mu_{i}^{K}(t,C_{\ell_{2}}^{K})\bigg{)}\left[\frac{(\ell_{1}+\ell_{2})^{2}-\ell_{1}^{2}-\ell_{2}^{2}}{s_{K}^{2}}\right]$
$=\frac{\alpha_{\mathrm{max}}}{s_{K}^{2}}\sum_{\ell_{1},\ell_{2}\in\mathbb{N}}\bigg{(}\sum_{i\in[d]}\mu_{i}^{K}(t,C_{\ell_{1}}^{K})\bigg{)}\bigg{(}\sum_{i\in[d]}\mu_{i}^{K}(t,C_{\ell_{2}}^{K})\bigg{)}\ell_{1}\ell_{2}.$
Recalling the definition of $\mu_{i}^{K}$ yields
$\sum_{i\in[d]}\sum_{\ell\in{\mathbb{N}}}\mu_{i}^{K}(t,C_{\ell}^{K})\,\ell=\frac{N_{K}}{K}.$
Using this and Fubini yields $A^{K}_{C}H^{F_{1},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K}(t))\leq b^{2}\,\alpha_{\mathrm{max}}.$ Thus, setting $\psi_{1}(t)\coloneqq\mathbb{E}\left[H^{F_{1},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K}(t))\right]$, it then follows from Dynkin’s formula that
$\frac{\mathrm{d}}{{\mathrm{d}t}}\psi_{1}(t)\leq b^{2}\alpha_{\mathrm{max}},$
Moreover, at time 00 we have
$\psi_{1}(0)=\frac{1}{K}{\mathbb{E}}\bigg{[}\sum_{i\in[d]}\sum_{j\in[L_{i}^{K}(0)]}\frac{1}{s_{K}^{2}}\bigg{]}=\frac{\gamma_{K}}{s_{K}^{2}}\leq\frac{b^{2}}{\gamma_{K}},$
and (3.4) follows.
For (3.5) we proceed in a similar way, but using the function $F_{2}({\boldsymbol{x}})\coloneq\big{(}\sum_{i\in[d]}x_{i}\big{)}^{2}$. Once more we have $A^{K}_{M}H^{F_{2},{\boldsymbol{f}}}(\boldsymbol{{\boldsymbol{p}}})=0$. For the coalescence part we have
$A^{K}_{C}H^{F_{2},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K}(t))$ $\leq\frac{\alpha_{\mathrm{max}}}{2K}\sum_{\ell_{1},\ell_{2}\in\mathbb{N}}\sum_{i,j\in[d]}\mu_{i}^{K}(t,C_{\ell_{1}}^{K})\mu_{j}^{K}(t,C_{\ell_{2}}^{K})\left[2\bigg{(}\sum_{k\in[d]}\langle\mu_{k}^{K}(t)\|\cdot\|_{1}^{2}\rangle\bigg{)}\frac{(\ell_{1}+\ell_{2})^{2}-\ell_{1}^{2}-\ell_{2}^{2}}{s_{K}^{2}}\right.$
$\qquad\,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left.+\frac{\big{(}(\ell_{1}+\ell_{2})^{2}-\ell_{1}^{2}-\ell_{2}^{2}\big{)}^{2}}{s_{K}^{4}}\right]$
$=2b^{2}\alpha_{\mathrm{max}}\sum_{i\in[d]}\langle\mu_{i}^{K}(t),\|\cdot\|_{1}^{2}\rangle+\frac{2\alpha_{\mathrm{max}}}{K}\bigg{(}\sum_{i\in[d]}\langle\mu_{i}^{K}(t),\|\cdot\|_{1}^{2}\rangle\bigg{)}^{2}.$
Thus, setting $\psi_{2}(t)\coloneqq\mathbb{E}\left[H^{F_{2},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K}(t))\right]$, and combining Dynkin’s formula with (3.4) yields
$\frac{\mathrm{d}}{{\mathrm{d}t}}\psi_{2}(t)\leq 2\alpha_{\mathrm{max}}b^{4}\Big{(}\frac{1}{\gamma_{K}}+\alpha_{\mathrm{max}}t\Big{)}+\frac{2\alpha_{\mathrm{max}}}{K}\psi_{2}(t).$
Moreover, at time 00, we have $\psi_{2}(0)=b^{4}/\gamma_{K}^{2}$. We conclude that
$\psi_{2}(t)$ $\leq e^{\frac{2\alpha_{\mathrm{max}}}{K}t}\left[\psi_{2}(0)+\int_{0}^{t}e^{-\frac{2\alpha_{\mathrm{max}}}{K}}\Big{(}\frac{2\alpha_{\mathrm{max}}b^{4}}{\gamma_{K}}+2\alpha_{\mathrm{max}}^{2}b^{4}s\Big{)}\mathrm{d}s\right],$
and (3.5) follows. ∎
3.3. Convergence of the generators
In this section, we will prove a uniform convergence result for the generator $A^{K}$. Since our final goal is to prove the convergence of the measures $({\boldsymbol{\mu}}^{K})_{K\geq 1}$ toward the solution of the coagulation equation (2.3) (resp. (2.5)), the generator of the limiting object should be given formally by
$\bar{A}G({\boldsymbol{p}})=\frac{{\rm d}}{{\rm d}t}G\circ u(\cdot,{\boldsymbol{p}})|_{t=0}$
where $u(\cdot,{\boldsymbol{p}})$ denotes the solution of Eq. (2.3) (resp. Eq. (2.5)) with initial value ${\boldsymbol{p}}$. For functions $G=H^{F,{\boldsymbol{f}}}\in{\mathcal{H}}$, a straightforward calculation shows that $\bar{A}H^{F,{\boldsymbol{f}}}({\boldsymbol{p}})=\bar{A}_{M}H^{F,{\boldsymbol{f}}}({\boldsymbol{p}})+\bar{A}_{C}H^{F,{\boldsymbol{f}}}({\boldsymbol{p}})\text{ with}$
$\bar{A}_{M}H^{F,{\boldsymbol{f}}}({\boldsymbol{p}})$ $=\sum_{i\neq j}w_{i,j}\left[\langle f_{j},p_{i}\rangle\cdot\partial_{j}F(\langle{\boldsymbol{p}},{\boldsymbol{f}}\rangle)-\langle f_{i},p_{i}\rangle\cdot\partial_{i}F(\langle{\boldsymbol{p}},{\boldsymbol{f}}\rangle)\right]$
$\bar{A}_{C}H^{F,{\boldsymbol{f}}}({\boldsymbol{p}})$ $=\frac{1}{2}\sum_{i\in[d]}\left(\int f_{i}({\boldsymbol{c}})(p_{i}\star p_{i})(d{\boldsymbol{c}})-2\langle 1,p_{i}\rangle\langle f_{i},p_{i}\rangle\right)\partial_{i}F(\langle{\boldsymbol{p}},{\boldsymbol{f}}\rangle)\text{.}$
The next result formalises the announced convergence result.
Proposition 3.2.
For any $F\in C_{2}({\mathbb{R}}^{d})$ with bounded second order partial derivatives and ${\boldsymbol{f}}\coloneqq(f_{i})_{i\in[d]}$ with $f_{i}\in C_{b}(\mathbb{R}_{+}^{d})$, for $i\in[d]$, there is a constant $C(F,{\boldsymbol{f}})>0$ such that
$|(A^{K}-\bar{A})H^{F,{\boldsymbol{f}}}({\boldsymbol{p}})|\leq\frac{C(F,{\boldsymbol{f}})}{K}\bigg{(}\sum_{i\in[d]}p_{i}([N_{K}]_{0}/s_{K})+p_{i}([N_{K}]_{0}/s_{K})^{2}\bigg{)},\quad{\boldsymbol{p}}\in{\mathcal{M}}_{f}([N_{K}]_{0}^{d}/s_{K})^{d}.$ (3.6)
Proof.
Let us start with the migration part of the generator. A second order Taylor’s expansion for $F$ at $\langle{\boldsymbol{p}},{\boldsymbol{f}}\rangle$ yields
$F\bigg{(}\langle{\boldsymbol{p}},{\boldsymbol{f}}\rangle+\frac{{\boldsymbol{e}}_{j}f_{j}(\frac{{\boldsymbol{c}}}{s_{K}})-{\boldsymbol{e}}_{i}f_{i}(\frac{{\boldsymbol{c}}}{s_{K}})}{K}\bigg{)}-F(\langle{\boldsymbol{p}},{\boldsymbol{f}}\rangle)$ $=\partial_{j}F(\langle{\boldsymbol{p}},{\boldsymbol{f}}\rangle)\frac{f_{j}(\frac{{\boldsymbol{c}}}{s_{K}})}{K}-\partial_{i}F(\langle{\boldsymbol{p}},{\boldsymbol{f}}\rangle)\frac{f_{i}(\frac{{\boldsymbol{c}}}{s_{K}})}{K}+R_{M}^{K}({\boldsymbol{p}},{\boldsymbol{c}}),$
where $|R_{M}^{K}({\boldsymbol{p}},{\boldsymbol{c}})|\leq C_{F,{\boldsymbol{f}}}/K^{2},$ for a constant $C_{F,{\boldsymbol{f}}}>0$ depending only on the supremum of ${\boldsymbol{f}}$ and the second order partial derivatives of $F$. Plugging this into Eq. (3.2) yields
$A^{K}_{M}H^{F,{\boldsymbol{f}}}({\boldsymbol{p}})=\bar{A}_{M}H^{F,{\boldsymbol{f}}}({\boldsymbol{p}})+\epsilon_{M}^{K}({\boldsymbol{p}}),\quad\textrm{with}\quad|\epsilon_{M}^{K}({\boldsymbol{p}})|\leq\frac{C_{F,{\boldsymbol{f}}}w_{\mathrm{max}}}{K}\sum_{i\in[d]}p_{i}([N_{K}]_{0}/s_{K}),$
and $w_{\mathrm{max}}\coloneqq\max_{i,j\in[d]}w_{i,j}$.
Similarly, for the coalescence part of the generator, we use a second order Taylor expansion at $\langle{\boldsymbol{p}},{\boldsymbol{f}}\rangle$, to obtain
$F\bigg{(}\langle{\boldsymbol{p}},{\boldsymbol{f}}\rangle+\frac{\Delta_{i}{\boldsymbol{f}}(\frac{({\boldsymbol{c}}_{1},{\boldsymbol{c}}_{2})}{s_{K}})}{K}\bigg{)}-F(\langle{\boldsymbol{p}},{\boldsymbol{f}}\rangle)$ $=\partial_{i}F(\langle{\boldsymbol{p}},{\boldsymbol{f}}\rangle)\frac{f_{i}(\frac{{\boldsymbol{c}}_{1}+{\boldsymbol{c}}_{2}}{s_{K}})-f_{i}(\frac{{\boldsymbol{c}}_{1}}{s_{K}})-f_{i}(\frac{{\boldsymbol{c}}_{2}}{s_{K}}))}{K}+R_{C}^{K}({\boldsymbol{p}},{\boldsymbol{c}}_{1},{\boldsymbol{c}}_{2}),$
$|R_{C}^{K}({\boldsymbol{p}},{\boldsymbol{c}}_{1},{\boldsymbol{c}}_{2})|\leq\tilde{C}_{F,{\boldsymbol{f}}}/K^{2},$ for a constant $\tilde{C}_{F,{\boldsymbol{f}}}>0$ depending only on the supremum of ${\boldsymbol{f}}$ and the second order derivatives of $F$. Plugging this into Eq. (3.1), we obtain
$A^{K}_{C}H^{F,{\boldsymbol{f}}}({\boldsymbol{p}})=\bar{A}_{C}H^{F,{\boldsymbol{f}}}({\boldsymbol{p}})+\epsilon_{C}^{K}({\boldsymbol{p}}),\quad\textrm{with}\quad|\epsilon_{C}^{K}({\boldsymbol{p}})|\leq\frac{\tilde{C}_{F,{\boldsymbol{f}}}\alpha_{\mathrm{max}}}{K}\sum_{i\in[d]}\left(p_{i}([N_{K}]_{0}/s_{K})\right)^{2}.$
The result follows combining the bounds for the error arising from the migration and from the coalescent parts. ∎
3.4. Comparison results
In this section we will provide a comparison result between the total number of blocks in the multi-type coalescent ${\boldsymbol{\Pi}}^{K}$ and two Kingman coalescents. This result will allow us to get a hand on the order of magnitude of the colony sizes.
3.4.1. The coupling
Let $L^{K}\coloneqq(L^{K}(t))_{t\geq 0}$ be the process that accounts for the total number of blocks in ${\boldsymbol{\Pi}}^{K}$, i.e.
$L^{K}(t)\coloneqq\sum_{i\in[d]}L_{i}^{K}(t),\quad t\geq 0.$
We start this section with the announced coupling result.
Lemma 3.3 (Coupling to Kingman).
There is a coupling between the vector-valued process $(L_{i}^{K})_{i\in[d]}$ and other two processes ${\hat{L}}^{K}$ and $\tilde{L}^{K}$ distributed as the block-counting processes of Kingman coalescents with merger rates per pair of blocks $\alpha_{\mathrm{max}}\coloneqq\max_{i\in[d]}\alpha_{i}$ and $\alpha_{\mathrm{min}}(d)\coloneqq\min_{i\in[d]}\alpha_{i}/d^{2}$, respectively, such that
$\hat{L}^{K}(t)\leq L^{K}(t)\leq\tilde{L}^{K}(t)\vee(d+1),\quad\text{for all }t\geq 0,$
and $\hat{L}^{K}(0)=L^{K}(0)=\tilde{L}^{K}(0)=N_{K}$.
Proof.
We first prove deterministic inequalities that will help us to compare the coalescence rates of the three processes. We claim that for any ${\boldsymbol{\ell}}\coloneqq(\ell_{i})_{i\in[d]}\in{\mathbb{N}}_{0}^{d}$ with $|{\boldsymbol{\ell}}|\geq d+1$,
$\alpha_{\mathrm{min}}(d)|{\boldsymbol{\ell}}|(|{\boldsymbol{\ell}}|-1)\leq\sum_{i\in[d]}\alpha_{i}\ell_{i}(\ell_{i}-1)\leq\alpha_{\mathrm{max}}|{\boldsymbol{\ell}}|(|{\boldsymbol{\ell}}|-1).$ (3.7)
Indeed, for $\ell\in{\mathbb{N}}_{0}^{d}$ with $|{\boldsymbol{\ell}}|-1\geq d$, Cauchy-Schwartz inequality yields
${\sum_{i\in[d]}\ell_{i}(\ell_{i}-1)}={\|{\boldsymbol{\ell}}\|_{2}^{2}-|{\boldsymbol{\ell}}|}\geq{\frac{|{\boldsymbol{\ell}}|^{2}}{d}-|{\boldsymbol{\ell}}|}=|{\boldsymbol{\ell}}|\left(\frac{|{\boldsymbol{\ell}}|}{d}-1\right)=|{\boldsymbol{\ell}}|(|{\boldsymbol{\ell}}|-1)\left(\frac{1}{d}-\frac{1-\frac{1}{d}}{|{\boldsymbol{\ell}}|-1}\right)\geq\frac{1}{d^{2}}|{\boldsymbol{\ell}}|(|{\boldsymbol{\ell}}|-1),$
and the claimed lower bound for $\sum_{i\in[d]}\alpha_{i}\ell_{i}(\ell_{i}-1)$ follows. The upper bound is a direct consequence of the inequality $\sum_{i\in[d]}|a_{i}|^{2}\leq(\sum_{i\in[d]}|a_{i}|)^{2}$.
Equipped with (3.7) we construct the announced coupling. We start $(L_{i}^{K})_{i\in[d]}$, ${\hat{L}}^{K}$ and $\tilde{L}^{K}$ such that $\hat{L}^{K}(0)=L^{K}(0)=\tilde{L}^{K}(0)=N_{K}$.
Assume that we have constructed them up to the $k$-th transition (transitions refer to the times at which at least one of the three processes makes a jump; we consider time 00 as the 00-th transition), which brings $(L_{i}^{K})_{i\in[d]}$, ${\hat{L}}^{K}$ and $\tilde{L}^{K}$ respectively to states $\ell_{*},{\boldsymbol{\ell}},\ell^{*}$ satisfying $\ell_{*}\leq|{\boldsymbol{\ell}}|\leq\ell^{*}$ and $|{\boldsymbol{\ell}}|\geq d+1$. The next transition is then constructed as follows. We define
$\rho({\boldsymbol{\ell}})\coloneqq K\sum_{i\in[d]}\sum_{j\neq i}\ell_{i}w_{i,j},\quad\hat{c}(\ell_{*})\coloneqq\alpha_{\mathrm{max}}\binom{\ell_{*}}{2},\quad c({\boldsymbol{\ell}})\coloneqq\sum_{i\in[d]}\alpha_{i}\binom{\ell_{i}}{2},\quad\tilde{c}(\ell^{*})\coloneqq\alpha_{\mathrm{min}}(d)\binom{\ell^{*}}{2},$
and $C(\ell_{*},{\boldsymbol{\ell}},\ell^{*})\coloneqq\max\{\hat{c}(\ell_{*}),c({\boldsymbol{\ell}}),\tilde{c}(\ell^{*})\}$. We refer to $\hat{c}(\ell_{*}),c({\boldsymbol{\ell}})$ and $\tilde{c}(\ell^{*})$ as coalescence rates of $\ell_{*}$, ${\boldsymbol{\ell}}$, $\ell^{*}$, respectively. Then after an exponential time with parameter $\lambda(\ell_{*},{\boldsymbol{\ell}},\ell^{*})\coloneqq\rho({\boldsymbol{\ell}})+\max\{\hat{c}(\ell_{*}),c({\boldsymbol{\ell}}),\tilde{c}(\ell^{*})\}$
  1. (1)
    with probability $\rho({\boldsymbol{\ell}})/\lambda(\ell_{*},{\boldsymbol{\ell}},\ell^{*})$ a migration event takes place (affecting only ${\boldsymbol{\ell}}$). More precisely,
    ${\boldsymbol{\ell}}\to{\boldsymbol{\ell}}+{\boldsymbol{e}}_{j}-{\boldsymbol{e}}_{i},\quad\textrm{with probability $K\ell_{i}w_{i,j}/\rho({\boldsymbol{\ell}})$},$
    $\ell_{*},\ell^{*}$ (and $|{\boldsymbol{\ell}}|$) remain unchanged.
  2. (2)
    with probability $C(\ell_{*},{\boldsymbol{\ell}},\ell^{*})/\lambda(\ell_{*},{\boldsymbol{\ell}},\ell^{*})$ a coalescence event takes place. The coalescence will always produce a transition among states $\ell_{*},{\boldsymbol{\ell}},\ell^{*}$ whose coalescence rate is equal to $C(\ell_{*},{\boldsymbol{\ell}},\ell^{*})$; the state associated with the second highest coalescence rate will only be modified with some probability and only then will the state with the lowest coalescence rate be affected with some probability.
    There are $6$ possible cases to consider that correspond to the different orderings of the coalescence rates of $\ell_{*},{\boldsymbol{\ell}},\ell^{*}$. We explain in detail one case; the others are analogous. We consider the case where $\hat{c}(\ell_{*})\geq c({\boldsymbol{\ell}})\geq\tilde{c}(\ell^{*})$. In this case, the transition
    $\ell_{*}\to\ell_{*}-1$
    takes place. In addition, with probability $c({\boldsymbol{\ell}})/\hat{c}(\ell_{*})$ the state ${\boldsymbol{\ell}}$ transitions. The transition affects only one of its coordinates; it is the $i$-th coordinate with probability $\alpha_{i}\ell_{i}(\ell_{i}-1)/2c({\boldsymbol{\ell}})$, in which case the transition
    ${\boldsymbol{\ell}}\to{\boldsymbol{\ell}}-{\boldsymbol{e}}_{i}$
    takes place. Finally, only if ${\boldsymbol{\ell}}$ changes, the transition $\tilde{c}(\ell^{*})/c({\boldsymbol{\ell}})$
    $\ell^{*}\to\ell^{*}-1$
    occurs with probability $\tilde{c}(\ell^{*})/c({\boldsymbol{\ell}})$ ; otherwise, $\ell^{*}$ remains unchanged.
As soon as $|{\boldsymbol{\ell}}|$ goes below $d+1$, we continue to carry out the coupling between $\hat{L}^{K}$ and $L^{K}$; the process $\tilde{L}^{K}$ can then be further constructed independently.
Note that, thanks to (3.7), if $\ell_{*}=|{\boldsymbol{\ell}}|$, then $\hat{c}(\ell_{*})\geq c({\boldsymbol{\ell}})$. Similarly, if $\ell^{*}=|{\boldsymbol{\ell}}|\geq d+1$, then $\tilde{c}(\ell_{*})\leq c({\boldsymbol{\ell}})$. This, together with the fact that the transitions only decrease the size of the affected states by one, implies that the transitions will preserve the ordering $\ell_{*}\leq|{\boldsymbol{\ell}}|\leq\ell^{*}$. The proof is achieved by noticing that the so-constructed processes have the desired distributions. ∎
3.4.2. Kingman bounds
The following result on Kingman’s coalescent will be helpful in many proofs. We provide here a short proof for completeness.
Lemma 3.4 (Moment bounds).
Let $(L_{\rho}^{K}(t))_{t\geq 0}$ be the block-counting process of a Kingman coalescent with merger rate $\rho$, which started with $N_{K}$ blocks at time 00. Then, for any $p\geq 1$, we have
${\mathbb{E}}[(L_{\rho}^{K}(t))^{p}]\leq{\bigg{(}\frac{1}{N_{K}^{1/p}}+\frac{\rho}{4p}t\bigg{)}}^{-p},\qquad t\geq 0.$
Proof.
Let ${\mathcal{A}}$ denote the generator of $L_{\rho}^{K}$ and $\psi(n)=n^{p}$. Clearly,
${\mathcal{A}}\psi(n)=\rho\frac{n(n-1)}{2}((n-1)^{p}-n^{p}),\qquad n\geq 1.$
Using that, for $n\geq 1$, $n^{p}-(n-1)^{p}\geq n^{p-1}$ and $n(n-1)\geq n^{2}/2$, we get
${\mathcal{A}}\psi(n)\leq-\frac{\rho}{4}\psi(n)^{1+1/p}.$
Combining this with Dynkin’s formula and Jensen inequality, we get
${\mathbb{E}}[\psi(L_{\rho}^{K}(t))]\leq\psi(N_{K})-\frac{\rho}{4}\int_{0}^{t}{\mathbb{E}}[\psi(L_{\rho}^{K}(s))^{1+1/p}]{\rm d}s\leq\psi(N_{K})-\frac{\rho}{4}\int_{0}^{t}{\mathbb{E}}[\psi(L_{\rho}^{K}(s))]^{1+1/p}{\rm d}s.$
Setting $v(t)={\mathbb{E}}[\psi(L_{\rho}^{K}(t))]$, the previous inequality implies that
$v^{\prime}(s)\leq-\frac{\rho}{4}v(s)^{1+1/p}.$
Dividing both sides by $v(s)^{1+1/p}$ and integrating both sides of the resulting inequality between 00 and $t$ and rearranging terms yields the result. ∎
4. Tightness and characterization of accumulation points
In this section we will prove the tightness of the sequence $\{{{\boldsymbol{\mu}}}^{K}\}_{K\geq 1}$ and characterize its accumulation points, paving the way to the desired convergence result.
4.1. Tightness
The next proposition states the tightness of $\{{\boldsymbol{\mu}}^{K}\}_{K\geq 1}$ and is the main result of this section.
Proposition 4.1.
Let $t_{0}=0$ in the critical sampling case and $t_{0}=\varepsilon>0$ (with $\varepsilon>0$ fix, but arbitrary) in the large sampling case. For any $T>0$, the sequence of measure-valued processes $\left\{({\boldsymbol{\mu}}^{K}(t))_{t\in[t_{0},T]}\right\}_{K\geq 1}$ is tight in $D([t_{0},T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d})^{d},w))$, where $(w)$ stands for the weak topology,
4.1.1. The proof strategy
To prove Proposition 4.1, we will show that the sequence of product measures
${\mu_{\otimes}^{K}}\coloneqq\bigotimes_{i\in[d]}\mu_{i}^{K},\qquad K\geq 1,$
is tight. The tightness of the sequence $\{{\boldsymbol{\mu}}^{K}\}_{K\geq 1}$ will then follow from the continuous mapping theorem (see [8, Thm. 2.7]).
With this in mind, we introduce the following notation. Let $\mathcal{F}$ denote the space of functions $f_{\otimes}\in C_{b}(\mathbb{R}_{+}^{d^{2}})$ of the form
$f_{\otimes}({\boldsymbol{x}}_{1},\dots,{\boldsymbol{x}}_{d})=\sum_{k\in[n]}\prod_{i\in[d]}f_{k,i}({\boldsymbol{x}}_{i}),\qquad{\boldsymbol{x}}_{1},\dots,{\boldsymbol{x}}_{d}\in\mathbb{R}_{+}^{d},$
for some $n\in{\mathbb{N}}$ and $f_{k,i}\in C_{b}(\mathbb{R}_{+}^{d})$ for $k\in[n]$, $i\in[d]$. According to the Stone-Weierstrass theorem, ${\mathcal{F}}$ is dense in $C_{0}(\mathbb{R}^{d^{2}})$ the set of continuous functions vanishing at infinity.
Our proof of the tightness of $\{{\mu_{\otimes}^{K}}\}_{K\geq 1}$ in $D([t_{0},T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d^{2}}),w))$ follows a standard line of argumentation (see e.g. [16, 42, 41]), which, in our setting, amounts to proving that the following three conditions are met.
(P1) For every function $f_{\otimes}\in\mathcal{F}$, the sequence $\{(\langle\mu_{\otimes}^{K}(t),f_{\otimes}\rangle)_{t\in[t_{0},T]}\}_{K\geq 1}$ is tight in $D([t_{0},T],\mathbb{R})$.
(P2) $\limsup_{K\rightarrow\infty}\mathbb{E}\left[\sup_{t\in[t_{0},T]}\langle\mu_{\otimes}^{K}(t),\|\cdot\|_{2}^{2}\rangle\right]<\infty$.
(P3) Any accumulation point of $\left\{\mu_{\otimes}^{K}\right\}_{K\geq 1}$ in $D([t_{0},T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d^{2}}),w))$ belongs to $C([t_{0},T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d^{2}}),w))$.
The strategy is analogous to the one used in [28, Thm. 7.4]), but they use a version of (P3) considering accumulation points in $D([t_{0},T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d^{2}}),v))$, where $(v)$ stands for the vague topology. The fact that these three conditions imply the tightness of $\{{\mu_{\otimes}^{K}}\}_{K\geq 1}$ follows from [42, Thm. 1.1.8] (see Remark 4.2 below). According to Roelly’s criterium (see [39]), condition (P1) already implies tightness in $D([t_{0},T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d^{2}}),v))$, where $(v)$ stands for the vague topology. Conditions (P2) and (P3) allow one to go from vague convergence to weak convergence (see [42, Lemma. 1.1.9]).
Remark 4.2.
Property (P2) slightly differs from the property stated in [42, Thm. 1.1.8], which reads as
$\lim_{k\rightarrow\infty}\limsup_{K\rightarrow\infty}\sup_{t\in[0,T]}\langle\mu_{\otimes}^{K}(t),\varphi_{k}(\|\cdot\|_{2})\rangle=0\text{, in probability,}$, (4.1)
where $\varphi_{k}$ is a function satisfying that $\mathds{1}_{\{\|{\boldsymbol{x}}\|_{2}\geq k\}}\leq\varphi_{k}(\|{\boldsymbol{x}}\|_{2})\leq\mathds{1}_{\{\|{\boldsymbol{x}}\|_{2}\geq k-1\}}.$ To see that (P2) is a stronger condition, note that Markov’s inequality implies that, for all $t\geq t_{0}$,
$\langle\mu_{\otimes}^{K}(t),\varphi_{k}(\|\cdot\|_{2})\rangle\leq\langle\mu_{\otimes}^{K}(t),\mathds{1}_{\{\|\cdot\|_{2}\geq k-1\}}\rangle=\langle\mu_{\otimes}^{K}(t),\mathds{1}_{\{\|\cdot\|_{2}^{2}\geq(k-1)^{2}\}}\rangle\leq\frac{1}{(k-1)^{2}}\langle\mu_{\otimes}^{K}(t),\|\cdot\|_{2}^{2}\rangle.$
In particular, we have
$\mathbb{E}\bigg{[}\sup_{t\in[t_{0},T]}\langle\mu_{\otimes}^{K}(t),\varphi_{k}(\|\cdot\|_{2})\rangle\bigg{]}\leq\frac{1}{(k-1)^{2}}\mathbb{E}\bigg{[}\sup_{t\in[t_{0},T]}\langle\mu_{\otimes}^{K}(t),\|\cdot\|_{2}^{2}\rangle\bigg{]}.$
Another application of Markov’s inequality shows that, under (P2), condition (4.1) holds.
Remark 4.3.
In view of our proof strategy, it will be useful to notice that (recall (3.1)), we have for $f_{\otimes}\in\mathcal{F}$ as above and $\pi_{[d]}({\boldsymbol{x}})\coloneqq\prod_{i\in[d]}x_{i}$
$\langle\mu_{\otimes}^{K}(t),f_{\otimes}\rangle=\sum_{k\in[n]}H^{\pi_{[d]},{\boldsymbol{f}}_{k}}(\boldsymbol{\mu}^{K}(t)).$ (4.2)
The proofs of (P1), (P2) and (P3) will be given in Lemmas 4.7,4.8 and 4.9, respectively.
4.1.2. A bound on the quadratic variations
In the next lemma, an important role is played by the functions ${\pi}_{I}:{\mathbb{R}}^{d}\to{\mathbb{R}}$, $I\subseteq[d]$, defined via ${\pi}_{I}({\boldsymbol{x}})=\prod_{j\in I}x_{j}.$
Lemma 4.4 (Quadratic variation).
Let $t_{0}=0$ in the critical sampling case and $t_{0}=\varepsilon>0$ in the large sampling case. For any ${\boldsymbol{f}}\coloneqq(f_{i})_{i\in[d]}$, $f_{i}\in C_{b}(\mathbb{R}_{+}^{d})$ and $I\subseteq[d]$, there is a constant $C({I},{\boldsymbol{f}})>0$ and decreasing processes $({\mathcal{I}}_{K}(s))_{s\geq t_{0}}$ and $({\mathcal{I}}_{K}^{*}(s))_{s\geq t_{0}}$ satisfying that
$\sup_{K\geq 1,s\geq t_{0}}{\mathbb{E}}[{\mathcal{I}}_{K}(s)]<\infty,\quad\sup_{K\geq 1,s\geq t_{0}}{\mathbb{E}}[{\mathcal{I}}_{K}^{*}(s)]<\infty,$
and such that for any $t\geq u\geq t_{0}$
$\left|\langle H^{{\pi}_{I},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K})\rangle_{t}-\langle H^{{\pi}_{I},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K})\rangle_{u}\right|$ $\leq\frac{C({I},{\boldsymbol{f}})}{K}{\mathcal{I}}_{K}(u)(t-u),$ (4.3)
$|A^{K}H^{{\pi}_{I},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K}(s))|$ $\leq{C}({I},{\boldsymbol{f}}){\mathcal{I}}_{K}^{*}(s),$ (4.4)
where $\langle\cdot\rangle_{t}$ stands for the quadratic variation at time $t$.
Proof.
We begin with the proof of (4.3). Recall the definition of $s_{K}$ from Eq. (2.1). We may then write
$\langle H^{\pi_{I},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K})\rangle_{t}-\langle H^{\pi_{I},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K})\rangle_{u}=\int_{u}^{t}I_{s}^{M}\mathrm{d}s+\int_{u}^{t}I_{s}^{C}\,\mathrm{d}s$
where
$I^{M}(s,K)\coloneqq\sum_{i\neq j}w_{i,j}\sum_{{\boldsymbol{c}}\in\mathbb{N}_{0}^{d}}K\mu^{K}_{i}\Big{(}s,\Big{\{}\frac{{\boldsymbol{c}}}{s_{K}}\Big{\}}\Big{)}\bigg{[}\pi_{I}\bigg{(}\langle\boldsymbol{\mu}^{K}(s),{\boldsymbol{f}}\rangle+\frac{\Delta_{i,j}{\boldsymbol{f}}(\frac{{\boldsymbol{c}}}{s_{K}})}{K}\bigg{)}-\pi_{I}\left(\langle\boldsymbol{\mu}^{K}(s),{\boldsymbol{f}}\rangle\right)\bigg{]}^{2},$
$I^{C}(s,K)\coloneqq\sum_{i\in[d]}\frac{\alpha_{i}K}{2}\sum_{{\boldsymbol{c}}_{1}\neq{\boldsymbol{c}}_{2}}\mu^{K}_{i}\Big{(}s,\Big{\{}\frac{{\boldsymbol{c}}_{1}}{s_{K}}\Big{\}}\Big{)}\mu^{K}_{i}\Big{(}s,\Big{\{}\frac{{\boldsymbol{c}}_{2}}{s_{K}}\Big{\}}\Big{)}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\bigg{[}\pi_{I}\bigg{(}\langle\boldsymbol{\mu}^{K}(s),{\boldsymbol{f}}\rangle+\frac{\Delta_{i}{\boldsymbol{f}}(\frac{{\boldsymbol{c}}_{1}}{s_{K}},\frac{{\boldsymbol{c}}_{2}}{s_{K}})}{K}\bigg{)}-\pi_{I}\left(\langle\boldsymbol{\mu}^{K}(s),{\boldsymbol{f}}\rangle\right)\bigg{]}^{2}$
$+\sum_{i\in[d]}\frac{\alpha_{i}}{2}\sum_{{\boldsymbol{c}}\in\mathbb{N}_{0}^{d}}\mu^{K}_{i}\Big{(}s,\Big{\{}\frac{{\boldsymbol{c}}}{s_{K}}\Big{\}}\Big{)}\Big{(}K\mu^{K}_{i}\Big{(}s,\Big{\{}\frac{{\boldsymbol{c}}}{s_{K}}\Big{\}}\Big{)}-1\Big{)}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\bigg{[}\pi_{I}\bigg{(}\langle\boldsymbol{\mu}^{K}(s),{\boldsymbol{f}}\rangle+\frac{\Delta_{i}{\boldsymbol{f}}(\frac{{\boldsymbol{c}}}{s_{K}},\frac{{\boldsymbol{c}}}{s_{K}})}{K}\bigg{)}-\pi_{I}\left(\langle\boldsymbol{\mu}^{K}(s),{\boldsymbol{f}}\rangle\right)\bigg{]}^{2}.$
Using that
$\pi_{I}$ $\bigg{(}\langle\boldsymbol{\mu}^{K}(s),{\boldsymbol{f}}\rangle+\frac{\Delta_{i,j}{\boldsymbol{f}}(\frac{{\boldsymbol{c}}}{s_{K}})}{K}\bigg{)}-\pi_{I}\left(\langle\boldsymbol{\mu}^{K}(s),{\boldsymbol{f}}\rangle\right)$
$\leq\max_{h\in I}\frac{\|f_{h}\|_{\infty}^{|I|}}{K}\bigg{(}\prod_{h\in I\setminus\{i\}}\langle\mu_{h}^{K}(s),1\rangle+\prod_{h\in I\setminus\{j\}}\langle\mu_{h}^{K}(s),1\rangle+\frac{1}{K}\prod_{h\in I\setminus\{i,j\}}\langle\mu_{h}^{K}(s),1\rangle\bigg{)}$
and
$\pi_{I}\bigg{(}\langle\boldsymbol{\mu}^{K}(s),{\boldsymbol{f}}\rangle+\frac{\Delta_{i}{\boldsymbol{f}}(\frac{{\boldsymbol{c}}_{1}}{s_{K}},\frac{{\boldsymbol{c}}_{2}}{s_{K}})}{K}\bigg{)}-\pi_{I}\left(\langle\boldsymbol{\mu}^{K}(s),{\boldsymbol{f}}\rangle\right)\leq\max_{h\in I}\frac{3\|f_{h}\|_{\infty}^{|I|}}{K}\prod_{h\in I\setminus\{i\}}\langle\mu_{h}^{K}(s),1\rangle,$
since $\langle\mu_{i}^{K}(s),1\rangle\leq\frac{L^{K}(\frac{s}{K})}{K}$, we see that
$|I^{M}(s,K)|$ $\leq\frac{C_{M}({I},{\boldsymbol{f}})}{K}\left(\left(\frac{L^{K}(\frac{s}{K})}{K}\right)^{2|I|-1}+\frac{1}{K}\left(\frac{L^{K}(\frac{s}{K})}{K}\right)^{(2|I|-3)\mathds{1}_{\{|I|>1\}}}\right)$
$\text{and}\quad|I^{C}(s,K)|$ $\leq\frac{C_{C}({I},{\boldsymbol{f}})}{K}\left(\frac{L^{K}(\frac{s}{K})}{K}\right)^{2|I|},$
for some constants $C_{M}(I,{\boldsymbol{f}}),C_{C}({I},{\boldsymbol{f}})>0$ depending only on the functions $\pi_{I},{\boldsymbol{f}}$. Since the moments of $L^{K}(s/K)/K$ are bounded due to Lemma 3.3 and Lemma 3.4, the Eq. (4.3) follows. We obtain Eq. (4.4) by repeating the same arguments without the squares. ∎
4.1.3. On the proof of (P1)
To prove (P1), we will use a classic result of Aldous and Rebolledo (see [1], [21]), which consists of two ingredients; the first is the content of the next lemma.
Lemma 4.5.
In the critical (resp. large) sampling case, for every fix $t\geq 0$ (resp. $t>0$), the sequence $\{\langle\mu_{\otimes}^{K}(t),f_{\otimes}\rangle\}_{K\geq 1}$ is tight.
Proof.
Since
$|\langle\mu_{\otimes}^{K}(t),f_{\otimes}\rangle|=\left|\sum_{k\in[n]}\prod_{i\in[d]}\langle\mu_{i}^{K}(t),f_{k,i}\rangle\right|\leq n\max_{k\in[n],j\in[d]}\lVert f_{k,j}\rVert\prod_{i\in[d]}\langle\mu_{i}^{K}(t),1\rangle.$
In the critical sampling case, we have $\langle\mu_{i}^{K}(t),1\rangle\leq\gamma_{K}$ and $\gamma_{K}$ is bounded, and hence, the result follows. In the large sampling case, we use that
$\prod_{i\in[d]}\langle\mu_{i}^{K}(t),1\rangle\leq\frac{(L^{K}(\frac{t}{K}))^{d}}{K^{d}}.$
Using this, Markov inequality, Lemma 3.3 and Lemma 3.4, we get for $t>0$
${\mathbb{P}}(|\langle\mu_{\otimes}^{K}(t),f_{\otimes}\rangle|>M)$ $\leq c_{1}\frac{{\mathbb{E}}[(L^{K}(\frac{t}{K}))^{d}]}{MK^{d}}\leq c_{1}\frac{{\mathbb{E}}[({\tilde{L}}^{K}(\frac{t}{K}){\vee(d+1)})^{d}]}{MK^{d}}\leq\frac{c_{2}}{t^{d}M},$
for some constants $c_{1},c_{2}>0$. The result follows. ∎
The second ingredient is provided by the following result.
Lemma 4.6.
Let $t_{0}=0$ in the critical sampling case and $t_{0}=\varepsilon>0$ in the large sampling case. Let ${\boldsymbol{f}}\coloneqq(f_{i})_{i\in[d]}$, $f_{i}\in C_{b}(\mathbb{R}_{+}^{d})$, and $I\subseteq[d]$. Define
$B_{t}^{K,I,{\boldsymbol{f}}}$ $\coloneqq H^{\pi_{I},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K}(t_{0}))+\int_{t_{0}}^{t}A^{K}H^{\pi_{I},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K}(s))\mathrm{d}s$
$\text{and}\quad M_{t}^{K,I,{\boldsymbol{f}}}$ $\coloneqq H^{\pi_{I},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K}(t))-H^{\pi_{I},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K}(t_{0}))-\int_{t_{0}}^{t}A^{K}H^{\pi_{I},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K}(s))\mathrm{d}s.$
Then, for any $\delta>0$ and any pair of stopping times $(\tau,\sigma)$ such that $t_{0}\leq\tau\leq\sigma\leq\tau+\delta\leq T$, there are constants $\tilde{c}_{1}({I},{\boldsymbol{f}},t_{0}),\tilde{c}_{2}({I},{\boldsymbol{f}},t_{0})>0$ such that
$\mathbb{E}[|M_{\sigma}^{K,I,{\boldsymbol{f}}}-M_{\tau}^{K,I,{\boldsymbol{f}}}|]\leq\tilde{c}_{1}({I},{\boldsymbol{f}},t_{0})\sqrt{\frac{\delta}{K}}\quad\text{and}\quad\mathbb{E}[|B_{\sigma}^{K,I,{\boldsymbol{f}}}-B_{\tau}^{K,I,{\boldsymbol{f}}}|]$ $\leq\tilde{c}_{2}(\pi_{I},{\boldsymbol{f}},t_{0})\delta.$
In particular, the two quantities are bounded from above by a function of $\delta$ that goes to 00 as $\delta\rightarrow 0$.
Proof.
Using Jensen’s inequality and the Martingale property, we get
$\mathbb{E}\left[\left|M_{\sigma}^{K,I,{\boldsymbol{f}}}-M_{\tau}^{K,I,{\boldsymbol{f}}}\right|\right]^{2}$ $\leq\mathbb{E}\left[\left(M_{\sigma}^{K,I,{\boldsymbol{f}}}-M_{\tau}^{K,I,{\boldsymbol{f}}}\right)^{2}\right]=\mathbb{E}\left[\left(M_{\sigma}^{K,I,{\boldsymbol{f}}}\right)^{2}-\left(M_{\tau}^{K,I,{\boldsymbol{f}}}\right)^{2}\right].$
In addition, according to Eq. (4.3) from Lemma 4.4, we find
$\mathbb{E}\left[\left(M_{\sigma}^{K,I,{\boldsymbol{f}}}\right)^{2}-\left(M_{\tau}^{K,I,{\boldsymbol{f}}}\right)^{2}\right]=$ $\mathbb{E}\left[\langle H^{\pi_{I},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K})\rangle_{\sigma}-\langle H^{\pi_{I},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K})\rangle_{\tau}\right]\leq\frac{C({I},{\boldsymbol{f}})}{K}\sup_{K\geq 1,s\geq t_{0}}\mathbb{E}[{\mathcal{I}}_{K}(s)]\delta.$
Summarizing, taking the square-root,
$\mathbb{E}[|M_{\sigma}^{K,I,{\boldsymbol{f}}}-M_{\tau}^{K,I,{\boldsymbol{f}}}|]$ $\leq\sqrt{\frac{C({I},{\boldsymbol{f}})}{K}\sup_{K\geq 1,s\geq t_{0}}\mathbb{E}[{\mathcal{I}}_{K}(s)]\delta}\leq\sqrt{C({I},{\boldsymbol{f}})\sup_{K\geq 1,s\geq t_{0}}\mathbb{E}[{\mathcal{I}}_{K}(s)]\delta}\underset{\delta\rightarrow 0}{\longrightarrow}0,$
uniformly in $K$, $\sigma$ and $\tau$. For the second expectation, we use Eq. (4.4) to find,
$\mathbb{E}[|B_{\sigma}^{K,I,{\boldsymbol{f}}}-B_{\tau}^{K,I,{\boldsymbol{f}}}|]$ $\leq\mathbb{E}\left[\int_{\tau}^{\sigma}|A^{K}H^{\pi_{I},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K}(s))|\mathrm{d}s\right]\leq{C}({I},{\boldsymbol{f}})\sup_{K\geq 1,s\geq t_{0}}\mathbb{E}[{\mathcal{I}}_{K}^{*}(s)]\delta\underset{\delta\rightarrow 0}{\longrightarrow}0,$
uniformly in $K$, $\sigma$ and $\tau$. ∎
Lemma 4.7 ((P1)).
Let $t_{0}=0$ in the critical sampling case and $t_{0}=\varepsilon>0$ in the large sampling case. For every function $f_{\otimes}\in\mathcal{F}$ and $t_{0}<T$, the sequence $\{(\langle{\mu_{\otimes}^{K}}(t),f_{\otimes}\rangle)_{t\in[0,T]}\}_{K\geq 1}$ is tight in $D([t_{0},T],\mathbb{R})$.
Proof.
The result follows as a consequence of Lemma 4.5 and Lemma 4.6 and the Aldous and Rebolledo criterium ([1], [21]). ∎
4.1.4. On the proof of (P2)
Now we proceed with the proof of condition (P2).
Lemma 4.8 ((P2)).
Let $t_{0}=0$ in the critical sampling case and $t_{0}=\varepsilon>0$ in the large sampling case. We have
$\limsup_{K\rightarrow\infty}\mathbb{E}\bigg{[}\sup_{t\in[t_{0},T]}\langle\mu_{\otimes}^{K}(t),\|\cdot\|_{2}^{2}\rangle\bigg{]}<\infty.$
Proof.
For the critical sampling, we would like to make use of Eq. (3.4) from Lemma 3.1. To that end, once again since $\langle\mu_{i}^{K}(t),1\rangle\leq b$, we find (with abuse of notation)
$\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\langle\mu_{\otimes}^{K}(t),\|\cdot\|_{2}^{2}\rangle\bigg{]}$ $=\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\sum_{i\in[d]}\langle\mu_{i}^{K}(t),\|\cdot\|_{2}^{2}\rangle\prod_{n\in[d]\setminus\{i\}}\langle\mu_{n}^{K}(t),1\rangle\bigg{]}\leq\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\sum_{i\in[d]}\langle\mu_{i}^{K}(t),\|\cdot\|_{2}^{2}\rangle\bigg{]}.$
Note that on the left hand side we integrate $\mu_{\otimes}^{K}(t)$ against the 2-Norm on $\mathbb{R}^{d^{2}}$, whereas on the right hand side we integrate $\mu_{i}^{K}(t)$ against the 2-Norm on $\mathbb{R}^{d}$. Further, since the sum of the squares is smaller than the square of the sum:
$\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\sum_{i\in[d]}\langle\mu_{i}^{K}(t),\|\cdot\|_{2}^{2}\rangle\bigg{]}\leq\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\sum_{i\in[d]}\langle\mu_{i}^{K}(t),\|\cdot\|_{1}^{2}\rangle\bigg{]}.$
Last, since the block sizes are only increasing over time, and again the sum of the squares is smaller than the square of the sum, Eq. (3.4) already yields
$\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\langle\mu_{\otimes}^{K}(t),\|\cdot\|_{2}^{2}\rangle\bigg{]}\leq\mathbb{E}\bigg{[}\sum_{i\in[d]}\langle\mu_{i}^{K}(T),\|\cdot\|_{1}^{2}\rangle\bigg{]}\leq b^{2}\Big{(}\frac{1}{\gamma_{K}}+T\Big{)},$
because we assumed $\alpha_{i}=1$ for all $i\in[d]$ throughout the section. Since $\gamma_{K}=N_{K}/K$, we conclude:
$\limsup_{K\rightarrow\infty}\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\langle\mu_{\otimes}^{K}(t),\|{\boldsymbol{x}}\|_{2}^{2}\rangle\bigg{]}\leq\limsup_{K\rightarrow\infty}b^{2}\Big{(}\frac{1}{\gamma_{K}}+T\Big{)}<\infty.$
The proof for the large sampling follows the same general idea, only that here, we make use of Eq. (3.5). More explicitly, this time we find (again with abuse of notation)
$\mathbb{E}\bigg{[}\sup_{t\in[\varepsilon,T]}\langle\mu_{\otimes}^{K}(t),\|\cdot\|_{2}^{2}\rangle\bigg{]}$ $=\mathbb{E}\bigg{[}\sup_{t\in[\varepsilon,T]}\sum_{i\in[d]}\langle\mu_{i}^{K}(t),\|\cdot\|_{2}^{2}\rangle\prod_{n\in[d]\setminus\{i\}}\langle\mu_{n}^{K}(t),1\rangle\bigg{]}$
$\leq\mathbb{E}\bigg{[}\sup_{t\in[\varepsilon,T]}\sum_{i\in[d]}\langle\mu_{i}^{K}(t),\|\cdot\|_{2}^{2}\rangle^{2}\bigg{]}+\mathbb{E}\bigg{[}\sup_{t\in[\varepsilon,T]}\sum_{i\in[d]}\bigg{(}\prod_{n\in[d]\setminus\{i\}}\langle\mu_{n}^{K}(t),1\rangle\bigg{)}^{2}\bigg{]}.$
Note that (again) on the left hand side we integrate $\mu_{\otimes}^{K}(t)$ against the 2-Norm on $\mathbb{R}^{d^{2}}$, whereas on the right hand side we integrate $\mu_{i}^{K}(t)$ against the 2-Norm on $\mathbb{R}^{d}$. For the first expectation on the right hand side, we again use that the sum of squares is smaller than the square of the sum and that blocks are only increasing over time:
$\mathbb{E}\bigg{[}\sup_{t\in[\varepsilon,T]}\sum_{i\in[d]}\langle\mu_{i}^{K}(t),\|\cdot\|_{2}^{2}\rangle^{2}\bigg{]}\leq\mathbb{E}\bigg{[}\bigg{(}\sum_{i\in[d]}\langle\mu_{i}^{K}(T),\|\cdot\|_{1}^{2}\rangle\bigg{)}^{2}\bigg{]}.$
We may therefore apply Eq. (3.5) to find
$\mathbb{E}\bigg{[}\sup_{t\in[\varepsilon,T]}\sum_{i\in[d]}\langle\mu_{i}^{K}(t),\|\cdot\|_{2}^{2}\rangle^{2}\bigg{]}\leq e^{\frac{2\alpha_{\mathrm{max}}}{K}T}b^{4}\Big{(}\frac{1}{\gamma_{K}^{2}}+\frac{2\alpha_{\mathrm{max}}}{\gamma_{K}}T+\alpha_{\mathrm{max}}^{2}T^{2}\Big{)}.$ (4.5)
For the second expectation, since the number of blocks is decreasing over time, we first find
$\mathbb{E}\bigg{[}\sup_{t\in[\varepsilon,T]}\sum_{i\in[d]}\bigg{(}\prod_{n\in[d]\setminus\{i\}}\langle\mu_{n}^{K}(t),1\rangle\bigg{)}^{2}\bigg{]}\leq d\mathbb{E}\bigg{[}\bigg{(}\frac{L^{K}(\frac{\varepsilon}{K})}{K}\bigg{)}^{2(d-1)}\bigg{]}.$
Moreover, according to Lemmas 3.3 and 3.4
$\mathbb{E}\bigg{[}\bigg{(}\frac{L^{K}(\frac{\varepsilon}{K})}{K}\bigg{)}^{2(d-1)}\bigg{]}\leq\frac{1}{K^{2(d-1)}}{\Bigg{(}\frac{1}{N_{K}^{\frac{1}{2(d-1)}}}+\frac{\rho_{0}}{8(d-1)}\frac{\varepsilon}{K}\Bigg{)}}^{-2(d-1)}.$ (4.6)
Combining Eq. (4.5) and Eq. (4.6), we arrive at
$\limsup_{K\rightarrow\infty}\mathbb{E}\bigg{[}\sup_{t\in[\varepsilon,T]}\langle\mu_{\otimes}^{K}(t),\|\cdot\|_{2}^{2}\rangle\bigg{]}\leq\alpha_{\mathrm{max}}^{2}T^{2}+\Big{(}\frac{8(d-1)}{\rho_{0}\varepsilon}\Big{)}^{2(d-1)}<\infty.\qed$
4.1.5. On the proofs of (P3) and Proposition 4.1
In this section we prove condition (P3) and conclude the proof of Proposition 4.1
Lemma 4.9 ((P3)).
Let $t_{0}=0$ in the critical sampling case and $t_{0}=\varepsilon>0$ in the large sampling case. Any accumulation point $\mu_{\otimes}^{\infty}$ of $\left\{\mu_{\otimes}^{K}\right\}_{K\geq 1}$ in $D([t_{0},T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d^{2}}),w))$ belongs to $C([t_{0},T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d^{2}}),w))$.
Proof.
We start with the critical sampling case. Let $\mu_{\otimes}^{\infty}$ be an accumulation point of the sequence $\left\{\mu_{\otimes}^{K}\right\}_{K\geq 1}$ in $D([0,T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d^{2}}),w))$. By a slight abuse of notation we denote by $\left\{\mu_{\otimes}^{K}\right\}_{K\geq 1}$ the subsequence converging to $\mu_{\otimes}^{\infty}$. Recall that
  • a migration from colony $j$ to colony $i$ moves a single point measure with a mass of $1/K$ from $\mu_{j}^{K}$ to $\mu_{i}^{K}$
  • a coalescence removes two point measures each with a mass of $1/K$ and adds another point measure, also with a mass of $1/K$.
Using this and that $\langle\mu_{i}^{K}(t),1\rangle\leq b$, it follows that
$\sup_{t\in[0,T]}$ $\sup_{f\in L^{\infty}(\mathbb{R}_{+}^{d^{2}}),\|f\|_{\infty}\leq 1}|\langle\mu_{\otimes}^{K}(t),f\rangle-\langle\mu_{\otimes}^{K}(t-),f\rangle|$
$\leq\sup_{t\in[0,T]}\frac{1}{K}\bigg{(}4\prod_{\ell\in[d]\setminus\{i\}}\langle\mu_{\ell}^{K}(t),1\rangle+\prod_{\ell\in[d]\setminus\{j\}}\langle\mu_{\ell}^{K}(t),1\rangle+\prod_{\ell\in[d]\setminus\{i,j\}}\frac{1}{K}\langle\mu_{\ell}^{K}(t),1\rangle\bigg{)}$
$\leq\frac{5b^{d-1}}{K}+\frac{b^{d-2}}{K^{2}}\underset{K\rightarrow\infty}{\longrightarrow}0\text{.}$
It follows that $\mu_{\otimes}^{\infty}\in C([0,T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d^{2}}),w))$.
Let us now consider the large sampling case. Similarly as before, let $\mu_{\otimes}^{\infty}$ be an accumulation point of the sequence $\left\{\mu_{\otimes}^{K}\right\}_{K\geq 1}$ in $D(\varepsilon,T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d^{2}}),w))$. By a slight abuse of notation we denote by $\left\{\mu_{\otimes}^{K}\right\}_{K\geq 1}$ the subsequence converging to $\mu_{\otimes}^{\infty}$. This time we obtain
$\sup_{t\in[\varepsilon,T]}$ $\sup_{f\in L^{\infty}(\mathbb{R}_{+}^{d^{2}}),\|f\|_{\infty}\leq 1}|\langle\mu_{\otimes}^{K}(t),f\rangle-\langle\mu_{\otimes}^{K}(t-),f\rangle|$
$\leq$ $\sup_{t\in[\varepsilon,T]}\frac{1}{K}\bigg{(}4\prod_{\ell\in[d]\setminus\{i\}}\langle\mu_{\ell}^{K}(t),1\rangle+\prod_{\ell\in[d]\setminus\{j\}}\langle\mu_{\ell}^{K}(t),1\rangle+\prod_{\ell\in[d]\setminus\{i,j\}}\frac{1}{K}\langle\mu_{\ell}^{K}(t),1\rangle\bigg{)}$
$\leq\frac{5}{K}\left(\frac{L^{K}(\frac{\varepsilon}{K})}{K}\right)^{d-1}+\frac{1}{K^{2}}\left(\frac{L^{K}(\frac{\varepsilon}{K})}{K}\right)^{d-2}\text{,}$
and the result follows using Lemma 3.3 and that, for the Kingman coalescent $\tilde{L}^{\infty}$ started with infinitely many lines, $t\tilde{L}_{t}^{\infty}$ converges almost surely as $t\to 0$ to a positive constant (see e.g. [5, Thm. 1] or [6, Chap. 2.1.2, Eq. (24)]). ∎
We now have all the ingredients to prove the tightness of $\{{\mu_{\otimes}^{K}}\}_{K\geq 1}$.
Proof of Proposition 4.1.
Thanks to Lemmas 4.7, 4.8 and 4.9 we conclude that conditions (P1), (P2) and (P3) are satisfied. Hence the result follows from [42, Thm. 1.1.8]. ∎
4.2. Characterization of the accumulation points
We conclude with the main result of this section, characterizing the accumulation points of the sequence of vectors of measures $\{\boldsymbol{\mu}^{K}\}_{K\geq 1}$:
Proposition 4.10.
Let $t_{0}=0$ in the critical sampling case and $t_{0}=\varepsilon>0$ in the large sampling case. Consider an accumulation point $\boldsymbol{\mu}^{\infty}$ of $\{\boldsymbol{\mu}^{K}\}_{K\geq 1}$. For every ${\boldsymbol{p}}\coloneqq(p_{i})_{i\in[d]}$ with $p_{i}\in D([t_{0},T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d}),w))$, we define
$\varphi_{t,{i},{\boldsymbol{f}}}({\boldsymbol{p}})$ $\coloneqq H^{F_{i},{\boldsymbol{f}}}({\boldsymbol{p}}(t))-H^{F_{i},{\boldsymbol{f}}}({\boldsymbol{p}}(t_{0}))-\int_{t_{0}}^{t}\bar{A}H^{F_{i},{\boldsymbol{f}}}({\boldsymbol{p}}(s))\mathrm{d}s$
[…]
for every $t\in[t_{0},T]$ and any $i\in[d]$.
Proof.
We prove the result for the large sampling case. By a slight abuse of notation we denote by $\{\boldsymbol{\mu}^{K}\}_{K\geq 1}$a subsequence converging to $\boldsymbol{\mu}^{\infty}$. We then observe from Proposition 3.2 and Lemma 4.4
$\mathbb{E}[|\varphi_{t,{i},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K})|]$ $\leq\mathbb{E}[|M_{t}^{K,\{i\},{\boldsymbol{f}}}-M_{0}^{K,\{i\},{\boldsymbol{f}}}|]+\mathbb{E}\left[\int_{\varepsilon}^{t}|(A^{K}-\bar{A})H^{F_{i},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K}(s))|\mathrm{d}s\right]$
The first expectation goes to 00 with $K\to\infty$ according to Lemma 4.6. For the second expectation, we find by Proposition 3.2
$\mathbb{E}\left[\int_{\varepsilon}^{t}|(A^{K}-\bar{A})H^{F_{i},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K}(s))|\mathrm{d}s\right]$ $\leq\mathbb{E}\bigg{[}\int_{\varepsilon}^{t}\frac{C(F_{i},{\boldsymbol{f}})}{K}\bigg{(}\sum_{i\in[d]}\langle\mu_{i}^{K}(t),1\rangle+\langle\mu_{i}^{K}(t),1\rangle^{2}\bigg{)}\mathrm{d}s\bigg{]}$
$\leq\frac{C(F_{i},{\boldsymbol{f}})(t-\varepsilon)}{K}\mathbb{E}\bigg{[}\frac{L^{K}(\frac{\varepsilon}{K})}{K}+\left(\frac{L^{K}(\frac{\varepsilon}{K})}{K}\right)^{2}\bigg{]}\underset{K\to\infty}{\longrightarrow}0,$
due to Lemma 3.4.
On the other hand, since any accumulation point $\boldsymbol{\mu}^{\infty}$ must be in $C([\varepsilon,T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d})^{d},w))$, because $\mu_{\otimes}^{\infty}\in C([\varepsilon,T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d^{2}}),w))$ and $\mu_{\otimes}^{\infty}\mapsto\boldsymbol{\mu}^{\infty}$ is continuous (and since for every continuous bounded $g$ the process $(\langle p^{K}(t),g\rangle)_{t\geq 0}$ converges to $(\langle p^{\infty}(t),g\rangle)_{t\geq 0}$ in the uniform norm on every finite interval if $(p^{K}(t))_{t\geq 0}$ converges to a continuous $(p^{\infty}(t))_{t\geq 0}$) we also have
$\varphi_{t,{i},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K})\Rightarrow\varphi_{t,{i},{\boldsymbol{f}}}(\boldsymbol{\mu}^{\infty})\text{.}$
Therefore we find
$\mathbb{E}[|\varphi_{t,{i},{\boldsymbol{f}}}(\boldsymbol{\mu}^{\infty})|]=\lim_{h\rightarrow\infty}\mathbb{E}[|\varphi_{t,{i},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K})|]=0$
by the uniform integrability of $\{|\varphi_{t,{i},{\boldsymbol{f}}}(\boldsymbol{\mu}^{K})|\}_{K\geq 1}$. Therefore, $\varphi_{t,{i},{\boldsymbol{f}}}(\boldsymbol{\mu}^{\infty})=0$ a.s for every $t\in[\varepsilon,T]$ and any $i\in[d]$. ∎
5. Proofs of Theorems  2.4 and 2.5.
In this section, we will provide proofs of our main results. We will start this section with the proof of Theorem  2.4, which will follow as a combination of the results obtained in the previous sections. The proof of Theorem 2.5 will require some additional work due to the degenerate initial condition.
5.1. The critical sampling case
As anticipated, we already have all the necessary ingredients to prove the main result in the critical sampling case.
Proof of Theorem 2.4.
We know from Proposition 4.1 that the sequence $\left\{\mu^{K}\right\}_{K\geq 1}$ is tight. Furthermore, Proposition 4.10 tells us that its accumulation points are weak solutions of Eq. (2.3). Moreover, under Assumption 2.3 we have
$\mu^{K}_{i}(0,\{{\boldsymbol{n}}\})=\frac{L^{K}_{i}(0)}{K}\delta_{e_{i},{\boldsymbol{n}}}\xrightarrow[K\to\infty]{}c\beta_{i}\delta_{e_{i},{\boldsymbol{n}}}.$ (5.1)
Hence, the result follows from the uniqueness of solutions to Eq. (2.3) with initial condition (2.4) (see Proposition A.1 in Appendix  A). ∎
5.2. The large sampling case
We start this section with the last ingredients for the proof of Theorem 2.5, namely the convergence (in an appropriate sense) of the initial conditions of the empirical measures $({\boldsymbol{\mu}}^{K})_{K\geq 1}$. Recall that the problem arises in the large sampling case because
$\langle\mu_{i}^{K}(0),1\rangle=\frac{L_{i}^{K}(0)}{K}\xrightarrow[K\to\infty]{}\infty.$
The next result allows us to circumvent this problem.
Proposition 5.1.
We have, for $\boldsymbol{\lambda}=(\lambda_{i})_{i\in[d]}>\boldsymbol{0}$,
$\lim_{\varepsilon\downarrow 0}\lim_{K\to\infty}\langle\mu_{i}^{K}(\varepsilon),1-e^{-\langle\boldsymbol{\lambda},\cdot\,\rangle}\rangle=\lambda_{i}\beta_{i},\textrm{ in probability.}$.
Before we dive into the proof of this result, let us use it to prove Theorem 2.5.
Proof of Theorem 2.5.
Let $\left\{{\boldsymbol{\mu}}^{K_{n}}\right\}_{n\geq 1}$ be an arbitrary subsequence of $\left\{{\boldsymbol{\mu}}^{K}\right\}_{K\geq 1}$. Due to Propositions 4.1 and 4.10, for each $\varepsilon\in(0,T)$, one can extract a subsequence $\{{\boldsymbol{\mu}}^{K_{n}^{{}^{\prime}}}\}_{n\geq 1}$ of $\left\{{\boldsymbol{\mu}}^{K_{n}}\right\}_{n\geq 1}$ converging in distribution to a measure-valued process $(\boldsymbol{u}^{\varepsilon}(t,\cdot))_{t\in[\varepsilon,T]}$ evolving according to (2.5). Then a standard diagonalization argument yields the existence of a further subsequence $\{{\boldsymbol{\mu}}^{K_{n}{{}^{\prime\prime}}}\}_{n\geq 1}$ of $\{{\boldsymbol{\mu}}^{K_{n}}\}_{n\geq 1}$ and a measure-valued process $(\boldsymbol{u}(t,\cdot))_{t\in(0,T]}$ such that
  • $\boldsymbol{u}$ evolves according to Eq. (2.5),
  • for any $\varepsilon\in(0,T)$, $\boldsymbol{u}|_{[\varepsilon,T]}=\boldsymbol{u}^{\varepsilon}$,
  • for each $\varepsilon\in(0,T)$, ${\boldsymbol{\mu}}^{K_{n}^{{}^{\prime\prime}}}\Rightarrow\boldsymbol{u}$ in $[\varepsilon,T]$.
We now claim that $\boldsymbol{u}$ for all $i\in[d]$,
[…]
If the claim is true, then by Theorem 2.7 the function $\boldsymbol{u}$ would be uniquely determined (it is then in fact a deterministic measure-valued function) and would not depend on the choice of the initial subsequence $\left\{{\boldsymbol{\mu}}^{K_{n}}\right\}_{n\geq 1}$. The result would then follow.
Let us now prove the claim. To simplify the notation, we will denote from now on the subsequence $\{\mu^{K_{n}{{}^{\prime\prime}}}\}_{n\geq 1}$ as the original sequence $\{\mu^{K}\}_{K\geq 1}$. By construction, we have for any $\varepsilon>0$
$\langle\mu^{K}_{i}(\varepsilon),1-e^{-\langle\boldsymbol{\lambda},\cdot\rangle}\rangle\Rightarrow\langle u_{i}(\varepsilon),1-e^{-\langle\boldsymbol{\lambda},\cdot\rangle}\rangle.$
Therefore, it follows that
$\liminf_{K\rightarrow\infty}\mathbb{P}(\langle\mu^{K}_{i}(\varepsilon),1-e^{-\langle\boldsymbol{\lambda},\cdot\rangle}\rangle\in G)\geq\mathbb{P}(\langle u_{i}(\varepsilon),1-e^{-\langle\boldsymbol{\lambda},\cdot\rangle}\rangle\in G),$
for every open set $G$. In particular, for any $\rho>0$:
$\mathbb{P}(|\langle\boldsymbol{u}_{i}(\varepsilon),1-e^{-\langle\boldsymbol{\lambda},\cdot\rangle}\rangle-\lambda_{i}\beta_{i}|>\rho)\leq\liminf_{K\rightarrow\infty}\mathbb{P}(|\langle\mu^{K}_{i}(\varepsilon),1-e^{-\langle\boldsymbol{\lambda},\cdot\rangle}\rangle-\lambda_{i}\beta_{i}|>\rho).$
[…]
$\limsup_{\varepsilon\downarrow 0}\mathbb{P}(|\langle u_{i}(\varepsilon),1-e^{-\langle\boldsymbol{\lambda},\cdot\rangle}\rangle-\lambda_{i}\beta_{i}|>\rho)$ $\leq\limsup_{\varepsilon\downarrow 0}\liminf_{K\rightarrow\infty}\mathbb{P}(|\langle\mu^{K}_{i}(\varepsilon),1-e^{-\langle\boldsymbol{\lambda},\cdot\rangle}\rangle-\lambda_{i}\beta_{i}|>\rho)=0.$
Hence $v_{i}(\varepsilon,{\boldsymbol{\lambda}})\coloneqq\langle u_{i}(\varepsilon),1-e^{-\langle\boldsymbol{\lambda},\cdot\rangle}\rangle$ converges in probability to $\lambda_{i}\beta_{i}$ as $\varepsilon\to 0$. We will now show that, for each $M\in{\mathbb{N}}$, the previous limit exists almost surely for all ${\boldsymbol{\lambda}}\geq 0$ (component-wise) with $\lVert{\boldsymbol{\lambda}}\rVert_{\infty}\leq M$, which combined with the convergence in probability would yield the desired result. Fix now $M\in{\mathbb{N}}$ and ${\boldsymbol{\lambda}}\geq 0$ with $\lVert{\boldsymbol{\lambda}}\rVert_{\infty}\leq M$.
Let $\varepsilon_{0}>0$ be arbitrary for the moment; we will choose it appropriately later. Define for $0<\varepsilon<\varepsilon_{0}$
$\theta_{i}^{\varepsilon}(t,\boldsymbol{\lambda})=\beta_{i}^{-1}\langle u_{i}(\varepsilon-t),1-e^{-\langle\boldsymbol{\lambda},\cdot\rangle}\rangle,$
and note that $\theta_{i}^{\varepsilon}(0,\boldsymbol{\lambda})=\beta_{i}^{-1}v_{i}(\varepsilon,{\boldsymbol{\lambda}})$. Since $\boldsymbol{u}$ solves (2.5), one can deduce that $(\theta_{i}^{\varepsilon})_{i\in[d]}$ solves the multi-dimensional ODE
$\partial_{t}\theta_{i}^{\varepsilon}\ =\ \frac{1}{2}(\alpha_{i}\beta_{i})(\theta_{i}^{\varepsilon})^{2}-\sum_{j\in[d]\setminus\{i\}}\bigg{(}w_{j,i}\frac{\beta_{j}}{\beta_{i}}\theta_{j}^{\varepsilon}-w_{i,j}\theta_{i}^{\varepsilon}\bigg{)},\ \mbox{ and $\theta_{i}^{\varepsilon}(0,\boldsymbol{\lambda})=\beta_{i}^{-1}v_{i}(\varepsilon,\boldsymbol{\lambda}).$}$
Note that $v_{i}(\varepsilon,\boldsymbol{\lambda})\leq v_{i}(\varepsilon,M\boldsymbol{1})$, where $\boldsymbol{1}$ is the $d$-dimensional vector with all coordinates equal to $1$. From the convergence in probability of $v_{i}(\varepsilon,M\boldsymbol{1})$ to $M\beta_{i}$, we infer that the event
$A_{M}(\varepsilon_{0})\coloneqq\bigcup_{\varepsilon<\varepsilon_{0}}\left\{v_{i}(\varepsilon,M\boldsymbol{1})\leq M\beta_{i}+\frac{1}{2}\right\}$
has probability $1$.
On the event $A_{M}(\varepsilon_{0})$, there is $\varepsilon_{*}\in(0,\varepsilon_{0})$ such that $v_{i}(\varepsilon_{*},\boldsymbol{\lambda})\leq v_{i}(\varepsilon_{*},M\boldsymbol{1})\leq M\beta_{i}+1/2$. Since, in addition, $\theta_{j}^{\varepsilon_{*}}\geq 0$ for all $j\in[d]$, it follows that $\theta_{i}^{\varepsilon_{*}}\leq\phi_{i}$, where $\phi_{i}$ solves the ODE
$\partial_{t}\phi_{i}\ =\ \frac{1}{2}(\alpha_{i}\beta_{i})\phi_{i}^{2}+w_{i}\phi_{i},\ \mbox{ and $\phi_{i}(0)=M+\frac{1}{2\beta_{i}},$}$
where $w_{i}=\sum_{j\in[d]\setminus\{i\}}w_{i,j}$. A simple explicit calculation shows that $\phi_{i}$ blows up at time
$t_{\infty}\coloneqq\frac{1}{w_{i}}\ln\left(1+\frac{2w_{i}}{\alpha_{i}(\beta_{i}M+\frac{1}{2})}\right)>0,$
and hence $\theta_{i}^{\varepsilon_{*}}$ does not blow up before $t_{\infty}$. Setting $\varepsilon_{0}\coloneqq t_{\infty}$, we conclude that, on $A_{M}(\varepsilon_{0})$, there is $\varepsilon_{*}\in(0,\varepsilon_{0})$ such that $\theta_{i}^{\varepsilon_{*}}(t,{\boldsymbol{\lambda}})$ admits a limit as $t\to\varepsilon_{*}$, which translates into the desired convergence of $\langle u_{i}(\varepsilon),1-e^{-\langle\boldsymbol{\lambda},\cdot\rangle}\rangle$ as $\varepsilon\to 0$. ∎
The proof of Proposition 5.1 builds on the following intuition (see Fig. 4). At small times, the coalescence rate is much higher than the migration rate. Therefore, we expect that, while coalescence reduces the number of blocks to $\mathcal{O}(K)$, the impact of migration is of a smaller order and can be neglected. More precisely, we will see that in a typical block at colony $i$ the number of elements of colors different from $i$ is much smaller than the scaling factor $\gamma_{K}$.
Figure 4. An illustration of the two scenarios. Left: mono-chromatic scenario (configurations at time $\varepsilon/K$: $(5,0,0)$, $(11,0,0)$, $(4,0,0)$ in colony $1$, $(0,5,0)$, $(0,7,0)$, $(0,4,0)$, $(0,7,0)$ in colony $2$, $(0,0,7)$, $(0,0,9)$, $(0,0,7)$ in colony $3$). Right: poly-chromatic scenario (configurations at time $\varepsilon/K$: $(11,0,0)$, $(0,6,0)$, $(0,2,7)$ in colony $1$, $(5,7,0)$, $(0,0,3)$ in colony $2$, $(4,8,0)$, $(0,0,13)$ in colony $3$).
To make this intuition precise, we will introduce two new measures $\bar{\mu}_{i}^{K}$ and $\Delta\mu_{i}^{K}$, the first accounting for the mono-chromatic part of the process and the second for the poly-chromatic part. To do this, we order the blocks in each colony according to their least element and denote by $B_{i,j,k}^{K}(t)$, $k\in[d]$, the number of elements of type $k$ in the $j$-th block in colony $i$. Note that ${\boldsymbol{B}}_{i,j}^{K}(t)\coloneqq(B_{i,j,k}^{K}(t))_{k\in[d]}$ is then the configuration of the $j$-th block in colony $i$. We then set
$\bar{\mu}_{i}^{K}(t)\coloneqq$ $\frac{1}{K}\sum_{j=1}^{L_{i}^{K}(\frac{t}{K})}\delta_{\frac{\bar{{\boldsymbol{B}}}_{i,j}^{K}(\frac{t}{K})}{\gamma_{K}}},\qquad\text{where }\bar{B}_{i,j,h}^{K}\Big{(}\frac{t}{K}\Big{)}=\begin{cases}B_{i,j,i}^{K}(\frac{t}{K}),\text{if }h=i\\ 0,\text{ otherwise,}\end{cases}$ (5.2)
$\Delta\mu_{i}^{K}(t)\coloneqq$ $\frac{1}{K}\sum_{j=1}^{L_{i}^{K}(\frac{t}{K})}\delta_{\frac{\Delta{\boldsymbol{B}}_{i,j}^{K}(\frac{t}{K})}{\gamma_{K}}},\qquad\text{where }\Delta B_{i,j,h}^{K}\Big{(}\frac{t}{K}\Big{)}=\begin{cases}0,\text{if }h=i\\ B_{i,j,h}^{K}(\frac{t}{K}),\text{otherwise.}\end{cases}$ (5.3)
Note that even if $\mu_{i}^{K}\neq\bar{\mu}_{i}^{K}+\Delta\mu_{i}^{K}$, we still have for $\boldsymbol{\lambda}\in\mathbb{R}_{+}^{d}$
$\mu_{i}^{K}(t)=\frac{1}{K}\sum_{j=1}^{L_{i}^{K}(\frac{t}{K})}\delta_{\frac{(\bar{{\boldsymbol{B}}}_{i,j}^{K}(\frac{t}{K})+\Delta{\boldsymbol{B}}_{i,j}^{K}(\frac{t}{K}))}{\gamma_{K}}},$
and hence
$\langle\mu_{i}^{K}(t),\langle\boldsymbol{\lambda},\cdot\,\rangle\rangle=\langle\bar{\mu}_{i}^{K}(t),\langle\boldsymbol{\lambda},\cdot\,\rangle\rangle+\langle\Delta\mu_{i}^{K}(t),\langle\boldsymbol{\lambda},\cdot\,\rangle\rangle.$
We will dedicate Subsections 5.2.1 and 5.2.2 respectively to the analysis of the monochromatic and the polychromatic parts in the previous expression.
5.2.1. Mono-chromatic part
The first thing to note here is that the total number of elements of color $i$ counted across all colonies remains constant, as migration simply moves elements from one colony to another. No elements change color, nor are they created or destroyed. So, if we only consider colony $i$, we lose elements of color $i$ due to migration to other colonies, although it is possible that some of them might return (also due to migration).
The next result, loosely speaking, states that we can neglect the impact of migration at small times $t$.
Lemma 5.2.
Let $\bar{\mu}_{i}^{K}$ be defined as in Eq. (5.2). Then
$\lim_{\varepsilon\downarrow 0}\lim_{K\to\infty}\mathbb{E}\left[\langle\bar{\mu}_{i}^{K}(\varepsilon),\langle\boldsymbol{\lambda},\cdot\rangle\rangle\right]=\lambda_{i}\beta_{i}.$
Proof.
Recall that we start with $L_{i}^{K}(0)$ singletons in colony $i$, and note that
$\langle\bar{\mu}_{i}^{K}(\varepsilon),\langle\boldsymbol{\lambda},\cdot\rangle\rangle=\frac{1}{K}\sum_{j=1}^{L_{i}^{K}(\frac{\varepsilon}{K})}\lambda_{i}\frac{\bar{B}_{i,j,i}^{K}\left(\frac{\varepsilon}{K}\right)}{\gamma_{K}}=\frac{\lambda_{i}}{N_{K}}\sum_{j=1}^{L_{i}^{K}(\frac{\varepsilon}{K})}\bar{B}_{i,j,i}^{K}\Big{(}\frac{\varepsilon}{K}\Big{)}=\lambda_{i}\frac{L_{i}^{K}(0)-E_{i}^{K}(\frac{\varepsilon}{K})}{N_{K}},$
where $E_{i}^{K}(s)$ denotes the number of elements of color $i$ that are not present in colony $i$ at time $s$. It remains to show
$\lim_{\varepsilon\downarrow 0}\lim_{K\rightarrow\infty}\frac{1}{N_{K}}\mathbb{E}\left[E_{i}^{K}\Big{(}\frac{\varepsilon}{K}\Big{)}\right]=0.$
To do this, we will construct a process $\hat{E}_{i}^{K}$ that upper bounds ${E}_{i}^{K}$; the key ingredients are
  1. (1)
    Migrations of blocks to colony $i$ without any element of color $i$ do not affect $E_{i}^{K}$. Thus, when constructing $\hat{E}_{i}^{K}$ we can remove at time 00 all singletons with a color different from $i$ and only keep the singletons of color $i$.
  2. (2)
    When a block in colony $i$ with $k$ elements of color $i$ migrates, it adds $k$ to $E_{i}^{K}$, but reduces the number of blocks in colony $i$, and therefore the rate of migration out of colony $i$. Thus, when constructing $\hat{E}_{i}^{K}$, at such a transition we will duplicate the block, keeping one copy in colony $i$ and the other contributing to $\hat{E}_{i}^{K}$. This way, when a block migrates back to colony $i$, its elements will have already duplicates in colony $i$, and hence can ignore that migration event. Note that in doing this, we are overcounting the number of elements of color $i$ out fo colony $i$, because the same element may contribute several times to $\hat{E}_{i}^{K}$, but not to $E_{i}^{K}$.
Having this in mind, we construct our upper-bound process $\hat{E}_{i}^{K}$ as follows (see Fig. 5).
  • We set $\hat{E}_{i}^{K}(0)=0$ and start a (single-type) Kingman coalescent with rate $\alpha_{i}$ with $L_{i}^{K}(0)$ singletons.
  • At rate $w_{i}K$, with $w_{i}\coloneqq\sum_{j\in[d]\setminus\{i\}}w_{i,j}$, per block of the Kingman coalescent, we count the number of elements within that block and increase by that amount the value of $\hat{E}_{i}^{K}$.
Figure 5. An illustration of $\hat{E}_{i}^{K}$. The circles on the vertical grey line on the right represent the times at which the number of elements within the blocks indicated by the arrows are counted.
Let $\hat{\mu}_{i}^{K}$ denote the rescaled empirical measure of the single-type Kingman coalescent described above. Then, for functions $F:\mathcal{M}_{f}(\mathbb{R}_{+})\times\mathbb{N}\rightarrow\mathbb{R}$, the generator of $\big{(}\hat{\mu}_{i}^{K}(t,\cdot),\hat{E}_{i}^{K}\left(t/K\right)\big{)}_{t\geq 0}$ is given as
$\hat{A}^{K}=\hat{A}^{K}_{M}+\hat{A}^{K}_{C}$
with
$\hat{A}^{K}_{M}F(p,m)=$ $w_{i}K\sum_{n\in\mathbb{N}}p\Big{(}\Big{\{}\frac{n}{\gamma_{K}}\Big{\}}\Big{)}\left[F\left(p,m+n\right)-F(p,m)\right],$
$\hat{A}^{K}_{C}F(p,m)=$ $\frac{\alpha_{i}K}{2}\sum_{n_{1}\neq n_{2}}p\Big{(}\Big{\{}\frac{n_{1}}{\gamma_{K}}\Big{\}}\Big{)}p\Big{(}\Big{\{}\frac{n_{2}}{\gamma_{K}}\Big{\}}\Big{)}\bigg{[}F\Big{(}p-\frac{\delta_{n_{1}+n_{2}}-\delta_{n_{1}}-\delta_{n_{2}}}{K},m\Big{)}-F(p,m)\bigg{]}$
$+\frac{\alpha_{i}K}{2}\sum_{n\in\mathbb{N}}p\Big{(}\Big{\{}\frac{n}{\gamma_{K}}\Big{\}}\Big{)}\Big{(}p\Big{(}\Big{\{}\frac{n}{\gamma_{K}}\Big{\}}\Big{)}-1\Big{)}\bigg{[}F\Big{(}p-\frac{\delta_{2n}-2\delta_{n}}{K},m\Big{)}-F(p,m)\bigg{]}.$
Consider first the function $F_{1}$ defined via $F_{1}(p,m)=\sum_{n\in\mathbb{N}}p\left(\left\{n/\gamma_{K}\right\}\right)n$. Note that $F_{1}$ is constant in its second component, and that it is invariant under coalescence events. Thus,
$\mathbb{E}\bigg{[}F_{1}\Big{(}\hat{\mu}_{i}^{K}(t,\cdot),\hat{E}_{i}\Big{(}\frac{t}{K}\Big{)}\Big{)}\bigg{]}=\mathbb{E}\bigg{[}F_{1}\Big{(}\hat{\mu}_{i}^{K}(0,\cdot),0\Big{)}\bigg{]}=\frac{L_{i}^{K}(0)}{K}.$ (5.4)
Consider now the function $F_{2}$ defined via $F_{2}(p,m)=m$. Since $F_{2}$ is constant in its first component, we have $\hat{A}^{K}_{C}F_{2}(p,m)=0$. Moreover, we have
$\hat{A}^{K}_{M}F_{2}(p,m)=w_{i}K\sum_{n\in\mathbb{N}}p\Big{(}\Big{\{}\frac{n}{\gamma_{K}}\Big{\}}\Big{)}n=w_{i}K\ F_{1}(p,m).$
Since $F_{2}(\hat{\mu}_{i}^{K}(0,\cdot),\hat{E}_{i}(0))=0$, using Dynkin’s formula and Eq. (5.4) yields
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}\Big{[}\hat{E}_{i}\Big{(}\frac{t}{K}\Big{)}\Big{]}=w_{i}K\ \mathbb{E}\bigg{[}F_{1}\Big{(}\hat{\mu}_{i}^{K}(t,\cdot),\hat{E}_{i}\Big{(}\frac{t}{K}\Big{)}\Big{)}\bigg{]}=w_{i}L_{i}^{K}(0),$
and therefore,
$\mathbb{E}\Big{[}\hat{E}_{i}\Big{(}\frac{t}{K}\Big{)}\Big{]}=w_{i}L_{i}^{K}(0)t.$
In particular
$0\leq\limsup_{K\rightarrow\infty}\frac{1}{N_{K}}\mathbb{E}\Big{[}E_{i}^{K}\Big{(}\frac{\varepsilon}{K}\Big{)}\Big{]}\leq\lim_{K\rightarrow\infty}\frac{1}{N_{K}}\mathbb{E}\Big{[}\hat{E}_{i}^{K}\Big{(}\frac{\varepsilon}{K}\Big{)}\Big{]}=w_{i}\beta_{i}\varepsilon.$
The result follows taking $\varepsilon\to 0$. ∎
5.2.2. Poly-chromatic part
The next result tells us that in colony $i$ the number of elements of a different color is much smaller than in $\gamma_{K}$.
Lemma 5.3.
Let $\Delta\mu_{i}^{K}$ be defined as in (5.3). Then
$\lim_{\varepsilon\downarrow 0}\lim_{K\to\infty}\mathbb{E}\left[\langle\Delta\mu_{i}^{K}(\varepsilon),\langle\boldsymbol{\lambda},\cdot\rangle\rangle\right]=0.$
Proof.
A straightforward calculation yields
$\langle\Delta\mu_{i}^{K}(\varepsilon),\langle\boldsymbol{\lambda},\cdot\rangle\rangle$ $=\frac{1}{K}\sum_{j=1}^{L_{i}^{K}(\frac{\varepsilon}{K})}\sum_{h\in[d]}\lambda_{h}\frac{\Delta B_{i,j,h}^{K}(\frac{\varepsilon}{K})}{\gamma_{K}}$
$\leq\frac{\lambda_{\mathrm{max}}}{N_{K}}\sum_{j=1}^{L_{i}^{K}(\frac{\varepsilon}{K})}\sum_{h\in[d]}\Delta B_{i,j,h}^{K}\Big{(}\frac{\varepsilon}{K}\Big{)}=\frac{\lambda_{\mathrm{max}}}{N_{K}}I_{i}^{K}\Big{(}\frac{\varepsilon}{K}\Big{)},$
where $\lambda_{\mathrm{max}}\coloneqq\max_{i\in[d]}\lambda_{i}$ and $I_{i}^{K}(t)$ denotes the number of elements, with a color different than $i$, present in colony $i$ at time $t$. It remains to show
$\lim_{\varepsilon\downarrow 0}\lim_{K\rightarrow\infty}\frac{1}{N_{K}}\mathbb{E}\Big{[}I_{i}^{K}\Big{(}\frac{\varepsilon}{K}\Big{)}\Big{]}=0.$
For this, note first that
$I_{i}^{K}(\varepsilon/K)\leq\sum_{j\in[d]\setminus\{i\}}{E}_{j}^{K}(\varepsilon/K),$
where ${E}_{j}^{K}$ is defined in the proof of Lemma 5.2. Using this and the upper bound $\hat{E}_{j}^{K}$ we obtained in the proof of Lemma 5.2, we get
$0\leq\limsup_{K\rightarrow\infty}\frac{1}{N_{K}}\mathbb{E}\Big{[}I_{i}^{K}\Big{(}\frac{\varepsilon}{K}\Big{)}\Big{]}$ $\leq\limsup_{K\rightarrow\infty}\frac{1}{N_{K}}\mathbb{E}\bigg{[}\sum_{j\in[d]\setminus\{i\}}\hat{E}_{j}^{K}\Big{(}\frac{\varepsilon}{K}\Big{)}\bigg{]}=\limsup_{K\rightarrow\infty}\frac{1}{N_{K}}\sum_{j\in[d]\setminus\{i\}}w_{j}L_{j}^{K}(0)\varepsilon$
$=\varepsilon\sum_{j\in[d]\setminus\{i\}}w_{j}\beta_{j}.$
The result follows taking $\varepsilon\to 0$. ∎
We conclude this section with the proof of Proposition 5.1.
Proof of Proposition 5.1.
Due to Markov’s inequality we have for all $\rho>0$:
$\mathbb{P}(|\langle\mu_{i}^{K}(t),1-e^{-\langle\boldsymbol{\lambda},\cdot\rangle}\rangle-\lambda_{i}\beta_{i}|>\rho)\leq\frac{1}{\rho}\mathbb{E}\big{[}\langle\mu_{i}^{K}(t),1-e^{-\langle\boldsymbol{\lambda},\cdot\rangle}\rangle-\lambda_{i}\beta_{i}\big{]}.$
Therefore, all that remains to do, is to show that
$\lim_{t\downarrow 0}\lim_{K\to\infty}\mathbb{E}\big{[}\langle\mu_{i}^{K}(t),1-e^{-\langle\boldsymbol{\lambda},\cdot\rangle}\rangle\big{]}=\lambda_{i}\beta_{i}.$
To accomplish this, we first write
$\langle\mu_{i}^{K}(\varepsilon),1-e^{-\langle\boldsymbol{\lambda},\cdot\rangle}\rangle=\langle\mu_{i}^{K}(\varepsilon),\langle\boldsymbol{\lambda},\cdot\rangle\rangle+R_{\boldsymbol{\lambda}}^{K}(\varepsilon).$
with $R_{\boldsymbol{\lambda}}^{K}(\varepsilon)=\langle\mu_{i}^{K}(\varepsilon),1-\langle\boldsymbol{\lambda},\cdot\rangle-e^{-\langle\boldsymbol{\lambda},\cdot\rangle}\rangle$. Since we have
$\langle\mu_{i}^{K}(\varepsilon),\langle\boldsymbol{\lambda},\cdot\rangle\rangle=\langle\bar{\mu}_{i}^{K}(\varepsilon),\langle\boldsymbol{\lambda},\cdot\rangle\rangle+\langle\Delta\mu_{i}^{K}(\varepsilon),\langle\boldsymbol{\lambda},\cdot\rangle\rangle,$
Lemmas 5.2 and 5.3 imply
$\lim_{\varepsilon\downarrow 0}\lim_{K\to\infty}\mathbb{E}\left[\langle\mu_{i}^{K}(\varepsilon),\langle\boldsymbol{\lambda},\cdot\rangle\rangle\right]=\lambda_{i}\beta_{i}.$
Moreover, since $x-x^{2}/2\leq 1-e^{-x}\leq x$, we infer that
$|R_{\boldsymbol{\lambda}}^{K}(\varepsilon)|\leq\frac{1}{2}\langle\mu_{i}^{K}(\varepsilon),\langle\boldsymbol{\lambda},\cdot\rangle^{2}\rangle\leq\frac{\lambda_{\mathrm{max}}^{2}}{2}\langle\mu_{i}^{K}(\varepsilon),\|\cdot\|_{1}^{2}\rangle\leq\frac{\lambda_{\mathrm{max}}^{2}}{2}\sum_{i\in[d]}\langle\mu_{i}^{K}(\varepsilon),\|\cdot\|_{1}^{2}\rangle.$
Therefore, using Eq. (3.4) from Lemma 3.1 we obtain
$\mathbb{E}[|R_{\boldsymbol{\lambda}}^{K}(\varepsilon)|]\leq\frac{\lambda_{\mathrm{max}}^{2}b^{2}}{2}\Big{(}\frac{1}{\gamma_{K}}+\alpha_{\mathrm{max}}\,\varepsilon\Big{)},$
and the result follows letting first $K\to\infty$ and then $\varepsilon\to 0$. ∎
Appendix A Uniqueness and stochastic representation for the critical sampling
Before diving into the proof of Theorem  2.8, let us prove the following result about the uniqueness of solutions to the discrete coagulation equation.
Proposition A.1.
The discrete coagulation equation (2.3) with initial condition (2.4) has at most one solution.
Proof.
Assume that ${\boldsymbol{u}}$ is a solution of (2.3) with initial condition (2.4). Define ${\boldsymbol{\rho}}\coloneqq\langle{\boldsymbol{u}},1\rangle$. A straightforward calculation shows that ${\boldsymbol{\rho}}$ solves the system of equations
$\partial_{t}\rho_{i}=-\frac{\alpha_{i}}{2}\rho_{i}^{2}+\sum_{j\in[d]\setminus\{i\}}w_{j,i}\rho_{j}-w_{i,j}\rho_{i},\quad i\in[d],$ (A.1)
with initial condition ${\boldsymbol{\rho}}(0)={c}{\boldsymbol{\beta}}$. According to Picard-Lindelöf theorem, that initial value problem has at most solution. Since we have one solution, that solution is uniquely determined.
Second, let us now fix ${\boldsymbol{n}}_{0}\in\mathbb{N}_{0}^{d}\setminus\{\boldsymbol{0}\}$. Note that ${\boldsymbol{u}}$ solves the system of non-linear odes
$\partial_{t}u_{i}(t,{\boldsymbol{n}})=\alpha_{i}\left(\frac{1}{2}\sum_{{\boldsymbol{n}}_{1}+{\boldsymbol{n}}_{2}={\boldsymbol{n}}}u_{i}(t,{\boldsymbol{n}}_{1})u_{i}(t,{\boldsymbol{n}}_{2})-\rho_{i}(t)u_{i}(t,{\boldsymbol{n}})\right)+\ \sum_{j}w_{ji}$ $u_{j}(t,{\boldsymbol{n}})-w_{ij}u_{i}(t,{\boldsymbol{n}}),$
$\qquad\qquad\quad i\in[d]\text{, }{\boldsymbol{n}}\in[{\boldsymbol{n}}_{0}],$
where ${\boldsymbol{\rho}}$ is the solution of the system of equations (A.1) with initial condition ${\boldsymbol{\rho}}(0)=c{\boldsymbol{\beta}}$. This is a finite dimensional system, and its solution is again uniquely determined by the initial condition, due to the Picard-Lindelöf theorem. The result follows as ${\boldsymbol{n}}_{0}$ was arbitrarily chosen. ∎
Now we proceed with the proof of Theorem 2.8 and therefore assume that ${\boldsymbol{\beta}}={\boldsymbol{\xi}}$ is the equilibrium probability measure, i.e.
$\sum_{j\in[d]\setminus\{i\}}\beta_{j}w_{j,i}\ =\ \beta_{i}\sum_{j\in[d]\setminus\{i\}}w_{i,j}.$
Proof of Theorem  2.8.
Let ${\boldsymbol{u}}$ be the solution of the discrete coagulation equation (2.3) under initial condition (2.4). For ${\boldsymbol{\lambda}}=(\lambda_{i})_{i\in[d]}\in[0,1]^{d}$, define $\hat{u}_{i}(t,{\boldsymbol{\lambda}})\coloneqq\sum_{{\boldsymbol{n}}\in\mathbb{N}_{0}^{d}\setminus\{\boldsymbol{0}\}}u_{i}(t,{\boldsymbol{n}}){\boldsymbol{\lambda}}^{\boldsymbol{n}}$ and
$\ v_{i}(t,{\boldsymbol{\lambda}})$ $=\left(1-\frac{\langle u_{i}(t),1\rangle}{{c}\beta_{i}}\right)\ +\ \frac{1}{{c}\beta_{i}}\hat{u}_{i}(t,{\boldsymbol{\lambda}}),$
Hence,
${c}\beta_{i}\partial_{t}v_{i}(t,\boldsymbol{\lambda})\ =\ -\partial_{t}\langle u_{i}(t),1\rangle+\sum_{{\boldsymbol{n}}\in\mathbb{N}_{0}^{d}\setminus\{\boldsymbol{0}\}}\partial_{t}\hat{u}_{i}(t,{\boldsymbol{n}}){\boldsymbol{\lambda}}^{\boldsymbol{n}}.$
For the first term we have
$\partial_{t}\langle u_{i},1\rangle=-\frac{\alpha_{i}}{2}\langle u_{i},1\rangle^{2}+\sum_{j\in[d]\setminus\{i\}}w_{j,i}\langle u_{j},1\rangle-w_{i,j}\langle u_{i},1\rangle.$ (A.2)
For the second term, recalling that ${\boldsymbol{u}}$ solves (2.3), we find
$\sum_{{\boldsymbol{n}}\in\mathbb{N}_{0}^{d}\setminus\{\boldsymbol{0}\}}\partial_{t}u_{i}(t,{\boldsymbol{n}}){\boldsymbol{\lambda}}^{\boldsymbol{n}}=$ $\alpha_{i}\bigg{(}\frac{1}{2}(\hat{u}_{i})^{2}-\langle u_{i},1\rangle\hat{u}_{i}\bigg{)}+\sum_{j\in[d]\setminus\{i\}}w_{j,i}\hat{u}_{j}-w_{i,j}\hat{u}_{i}$
$=$ $\alpha_{i}\bigg{(}\frac{{c^{2}}\beta_{i}^{2}}{2}\Big{(}v_{i}-\big{(}1-\frac{1}{{c}\beta_{i}}\langle u_{i},1\rangle\big{)}\Big{)}^{2}-\langle u_{i},1\rangle({c}\beta_{i}v_{i}-({c}\beta_{i}-|u_{i}|))\bigg{)}$
$+\sum_{j\in[d]\setminus\{i\}}w_{j,i}({c}\beta_{j}v_{j}-({c}\beta_{j}-\langle u_{j},1\rangle)-w_{i,j}({c}\beta_{i}v_{i}-({c}\beta_{i}-\langle u_{i},1\rangle).$ (A.3)
Combining the previous identities yields after a tedious, but straightforward calculation
$\partial_{t}v_{i}(t,{\boldsymbol{\lambda}})=$ $c\alpha_{i}\beta_{i}\left(\frac{1}{2}v_{i}^{2}+\frac{1}{2}-v_{i}\right)\ +\ \sum_{j\in[d]\setminus\{i\}}\frac{\beta_{j}}{\beta_{i}}v_{j}w_{j,i}-v_{i}w_{i,j}-\sum_{j\in[d]\setminus\{i\}}\frac{\beta_{j}}{\beta_{i}}w_{j,i}-w_{i,j}.$ (A.4)
In particular, if we set
$d_{i}\coloneqq\frac{c\alpha_{i}\beta_{i}}{2}-\sum_{j\in[d]\setminus\{i\}}\left(\frac{\beta_{j}}{\beta_{i}}w_{j,i}-w_{i,j}\right),$
we may rewrite (A.4) as
$\partial_{t}v_{i}(t,{\boldsymbol{\lambda}})=\ \frac{c\alpha_{i}\beta_{i}}{2}v_{i}^{2}+d_{i}-\bigg{(}\frac{c\alpha_{i}\beta_{i}}{2}+d_{i}\bigg{)}v_{i}+\sum_{j\in[d]\setminus\{i\}}\frac{\beta_{j}}{\beta_{i}}w_{j,i}(v_{j}-v_{i}).$ (A.5)
Recall that by assumption, for each $i\in[d]$, $d_{i}\geq 0$. Let us now consider a continuous-time multi-type branching process ${\boldsymbol{Z}}(t)=(Z_{i}(t))_{i\in[d]}$ such that, each particle of type $i\in[d]$
  • branches at rate $\frac{{c}\alpha_{i}\beta_{i}}{2}$,
  • dies at rate $d_{i}$,
  • makes a transition from $i$ to $j$ at rate $\frac{\beta_{j}}{\beta_{i}}w_{j,i}$.
It is straightforward to check that the function $\boldsymbol{h}(t,{\boldsymbol{\lambda}})\coloneqq{\mathbb{E}}_{{\boldsymbol{e}}_{i}}\left({\boldsymbol{\lambda}}^{{\boldsymbol{Z}}(t)}\right)$ satisfies the system of equations (A.5) and the initial condition $\boldsymbol{h}(0,\boldsymbol{\lambda})=\boldsymbol{\lambda}$ (see e.g. [3, Chap. V.7, Eq. (2)]). Thus, by uniqueness of this initial value problem, we infer that
$v_{i}(t,{\boldsymbol{\lambda}})\ =\ {\mathbb{E}}_{{\boldsymbol{e}}_{i}}\left({\boldsymbol{\lambda}}^{{\boldsymbol{Z}}(t)}\right),$
and the result follows. ∎
Remark A.2 (Stochastic representation at equilibrium).
Instead of assuming that $d_{i}\geq 0$ for all $i\in[d]$, let us assume that ${\boldsymbol{\beta}}={\boldsymbol{\xi}}$ is the equilibrium probability measure, i.e.
$\sum_{j\in[d]\setminus\{i\}}\beta_{j}w_{j,i}\ =\ \beta_{i}\sum_{j\in[d]\setminus\{i\}}w_{i,j}.$
We can then rewrite Eq. (A.4) as
$\partial_{t}v_{i}(t,{\boldsymbol{\lambda}})=$ $c\alpha_{i}\beta_{i}\left(\frac{1}{2}v_{i}^{2}+\frac{1}{2}-v_{i}\right)\ +\ \sum_{j\in[d]\setminus\{i\}}\frac{\beta_{j}}{\beta_{i}}v_{j}w_{j,i}-v_{i}w_{i,j}.$
Hence, proceeding as in the proof of Theorem 2.8, we deduce that
$u_{i}(t,{\boldsymbol{n}})\ =\ {c}\beta_{i}{\mathbb{P}}_{{\boldsymbol{e}}_{i}}\big{(}{\boldsymbol{Z}}(t)\ =\ {\boldsymbol{n}}\big{)},\quad t\geq 0,\,{\boldsymbol{n}}\in\mathbb{N}_{0}^{d}\setminus\{\boldsymbol{0}\},\,i\in[d],$
where ${\boldsymbol{Z}}(t)=(Z_{i}(t))_{i\in[d]}$ denotes a continuous-time multi-type branching process, such that a particle of color $i\in[d]$
  • branches and dies at rate $\frac{{c}\alpha_{i}\beta_{i}}{2}$.
  • makes a transition from $i$ to $j$ at rate $\frac{\beta_{j}}{\beta_{i}}w_{j,i}$.
Acknowledgements
The authors would like to thank Sebastian Hummel for insightful discussions at an early stage of the project. Fernando Cordero and Sophia-Marie Mellis are funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) — Project-ID 317210226 — SFB 1283.